### v1

# Multicut Problems in Almost-Planar Graphs: The Dependency of Complexity on the Demand Pattern

Abstract
Given a graph $G$, a set $T$ of terminal vertices, and a demand graph $H$ on $T$, the Multicut problem asks for a set of edges of minimum weight that separates the pairs of terminals specified by the edges of $H$. The Multicut problem can be solved in polynomial time if the number of terminals and the genus of the graph is bounded (Colin de Verdière [Algorithmica, 2017]). Restricting the possible demand graphs in the input leads to special cases of Multicut whose complexity might be different from the general problem.
Focke et al. [SoCG 2024] systematically characterized which special cases of Multicut are fixed-parameter tractable parameterized by the number of terminals on planar graphs. Moreover, extending these results beyond planar graphs, they precisely determined how the parameter genus influences the complexity and presented partial results of this form for graphs that can be made planar by the deletion of $\pi$ edges. Continuing this line of work, we complete the picture on how this parameter $\pi$ influences the complexity of different special cases and precisely determine the influence of the crossing number, another paramter measuring closeness to planarity.
Formally, let $\mathcal{H}$ be any class of graphs (satisfying a mild closure property) and let $\textsc{Multicut}(\mathcal{H})$ be the special case when the demand graph $H$ is in $\mathcal{H}$. Our fist main results is showing that if $\mathcal{H}$ has the combinatorial property of having bounded distance to extended bicliques, then $\textsc{Multicut}(\mathcal{H})$ on unweighted graphs is FPT parameterized by the number $t$ of terminals and $\pi$. For the case when $\mathcal{H}$ does not have this combinatorial property, Focke et al. [SoCG 2024] showed that $O(\sqrt{t})$ is essentially the best possible exponent of the running time; together with our result, this gives a complete understanding of how the parameter $\pi$ influences complexity on unweighted graphs. Our second main result is giving an algorithm whose existence shows that that the parameter crossing number behaves analogously if we consider $\textsc{Multicut}(\mathcal{H})$ on weighted graphs.
keywords:
MultiCut, Multiway Cut, Parameterized Complexity, Tight Bounds, Embedded Graph, Planar Graph, Crossing Number
1 Introduction
Most NP-hard problems remain NP-hard when restricted to planar graphs, unless they become trivial or irrelevant on planar graphs. There are very few NP-hard problems that are polynomial-time solvable on planar graphs for some nontrivial and interesting reason. One such case is the Multiway Cut problem: given a graph $G$ and a set $T\subseteq V(G)$ of terminals, the task is to find a multiway cut $S$ of minimum weight, that is, a set $S\subseteq E(G)$ such that every component of $G\setminus S$ contains at most one terminal. Dahlhaus et al. [5] showed that Multiway Cut in general graphs is $\mathsf{NP}$-hard even for three terminals, while it is polynomial-time solvable on planar graphs for every fixed number $|T|$ of terminals. The exponent of $n$ in the algorithm of Dahlhaus et al. [5] is $O(|T|)$. Later, the dependence was improved to $O(\sqrt{|T|})$ [13, 3], which was shown to be optimal under the Exponential-Time Hypothesis (ETH) [14, 2]. The results were extended to bounded-genus graphs, again with an essentially best possible form of the exponent [2, 3].
Multicut is a generalization of Multiway Cut where not all terminals need to be separated from each other, but the input specifies which pairs of terminals need to be separated (and we do not care whether other pairs are separated or not).
Multicut
Input: A weighted graph $G$ together with a set $T\subseteq V(G)$ of terminals;
and a demand graph $H$ with $V(H)=T$.
Output: A minimum-weight set $S\subseteq E(G)$ such that $u$ and $v$ are in distinct components
of $G\setminus S$ whenever $uv$ is an edge of $H$.
Let us observe that Multiway Cut is the special case of Multicut when $H$ is a complete graph on $T$. Colin de Verdière [3] presented an algorithm for Multicut where the exponent of the running time is $O(\sqrt{|T|})$, similarly to Multiway Cut. However, it is no longer clear anymore that this algorithm is a tight, optimal result and cannot be improved in some cases: it could be that there are special cases, special type of demand graphs where significantly better algorithms are possible. For example, if $H$ is a biclique (complete bipartite graph), then Multicut can be solved in polynomial time by a reduction to a minimum weight $s-t$ cut problem. A very natural question to consider is the case when $H$ is a complete 3-partite graph. This special case is the Group 3-Terminal Cut problem: given a graph $G$ with three sets of terminals $T_{1}$, $T_{2}$, $T_{3}$, the task is to find a set $S$ of edges of minimum total weight such that $G\setminus S$ has no path between any vertex of $T_{i}$ and any vertex of $T_{j}$ for all distinct $i,j\in[3]$. While a general Multicut algorithm solves Group 3-Terminal Cut with the exponent of the running time being the square root of the number of terminals, the lower bounds for Multiway Cut do not imply that this dependence is optimal.
Focke et al. [7] systematically analyzed which types of demand graphs make the problem more tractable. The question was explored in the following setting. Let $\mathcal{H}$ be a class of graphs (i.e., a set of graphs closed under isomorphism). Then we denote by $\mathcal{H}$) the special case of Multicut that contains only instances $(G,H)$ where $H\in\mathcal{H}$. For example, if $\mathcal{H}$ is the class of cliques, then $\mathcal{H}$) is exactly Multiway Cut.
Ideally, for every class $\mathcal{H}$ of graphs, we would like to understand the best possible running time of $\mathcal{H}$) on planar. In order to make the question more robust, Focke et al. [7] considered only classes $\mathcal{H}$ satisfying a mild closure property. We say that $H^{\prime}$ is a projection of $H$ if $H^{\prime}$ can be obtained from $H$ by deleting vertices and identifying pairs of independent vertices. Moreover, a class of graphs $\mathcal{H}$ is projection-closed if, whenever $H\in\mathcal{H}$ and $H^{\prime}$ is a projection of $H$, then $H^{\prime}\in\mathcal{H}$ also holds. The intuition is that an instance having demand graph $H^{\prime}$ can be easily simulated by an instance having demand graph $H$, thus it is reasonable to assume we only consider classes of demand graphs closed under this operation.
Let us consider some easy cases first. As we mentioned before, the Multicut problem can be solved in polynomial time by a reduction to $s-t$ minimum cut if the demand graph is a biclique. Furthermore, if $H$ has only two edges, then a similar reduction is known [10, 16]. Isolated vertices of $H$ clearly do not play any role in the problem. We say that $H$ is a trivial pattern if, after removing isolated vertices, it is either a biclique or has at most two edges. An extended biclique is a graph that consists of a complete bipartite graph together with a set of isolated vertices. Let $\mu$ be the minimum number of vertices that need to be deleted to make a graph $H$ an extended biclique. Then we say that $H$ has distance $\mu$ to extended bicliques. Moreover, we say that a graph class $\mathcal{H}$ has bounded distance to extended bicliques if there is a constant $\mu$ such that every $H\in\mathcal{H}$ has distance at most $\mu$ to extended bicliques.
For planar graphs, Focke et al. [7] give the following characterization:
Theorem 1.1 ([7]).
Let $\mathcal{H}$ be a computable projection-closed class of graphs. Then the following holds for $\mathcal{H}$).
  1. 1.
    If $\mathcal{H}$ has bounded distance to extended bicliques, then there is an $f(t)n^{O(1)}$ time algorithm.
  2. 2.
    Otherwise,
    1. (a)
      There is an $f(t)n^{O(\sqrt{t})}$ time algorithm.
    2. (b)
      Assuming ETH, there is a universal constant $\alpha>0$ such that for any fixed choice of $t\geq 3$, there is no $O(n^{\alpha\sqrt{t}})$ algorithm, even when restricted to unweighted instances with at most $t$ terminals.
That is, depending on whether $\mathcal{H}$ has bounded distance to extended bicliques, the number $t$ of terminals has to appear in the exponent. Focke et al. [7] precisely characterized how the complexity changes if we move from planar graphs to bounded-genus graphs. As we can see, the parameter genus can appear in the exponent in different ways, depending on $\mathcal{H}$.
Theorem 1.2 ([7]).
Let $\mathcal{H}$ be a computable projection-closed class of graphs. Then the following holds for $\mathcal{H}$).
  1. 1.
    If every graph in $\mathcal{H}$ is a trivial pattern, then there is a polynomial time algorithm.
  2. 2.
    Otherwise, if $\mathcal{H}$ has bounded distance to extended bicliques, then
    1. (a)
      There is an $f(g,t)n^{O(g)}$ time algorithm.
    2. (b)
      Assuming ETH, there is a universal constant $\alpha>0$ such that for any fixed choice of $\vec{g}\geq 0$, there is no $O(n^{\alpha(\vec{g}+1)/\log(\vec{g}+2)})$ algorithm, even when restricted to unweighted instances with orientable genus at most $\vec{g}$ and $t=3$ terminals.
  3. 3.
    Otherwise,
    1. (a)
      There is an $f(g,t)n^{O(\sqrt{g^{2}+gt+t})}$ time algorithm.
    2. (b)
      Assuming ETH, there is a universal constant $\alpha>0$ such that for any fixed choice of $\vec{g}\geq 0$ and $t\geq 3$, there is no $O(n^{\alpha\sqrt{\vec{g}^{2}+\vec{g}t+t}/\log(\vec{g}+t)})$ algorithm, even when restricted to unweighted instances with orientable genus at most $\vec{g}$ and at most $t$ terminals.
Focke et al. [7] also considered a much more restricted extension of planar graphs than bounded-genus graphs: graphs that can be made planar by deleting a few edges.We use $\textup{\sf planar}+\pi e$ for the class of graphs that can be made planar by removing at most $\pi$ edges. Given an instance $(G,H)$, we use $\pi$ for the smallest integer such that $G$ is in $\textup{\sf planar}+\pi e$.
Observe that in Theorem 1.2, the algorithmic results are stated for the edge-weighted version, while the lower bounds are stated for the unweighted version. In fact, edge-weights do not make much difference in Theorem 1.2, as polynomial edge weights can be simulated by parallel edges without changing the genus of the graphs. However, this simulation can easily change the number $\pi$ of edges that needs to be deleted to make the graph planar. This means that if we consider the influence of the parameter $\pi$ on the running time, the edge-weighted and the unweighted problem may behave differently. Focke et al. [7] characterized the influence of this parameter in the edge-weighted case:
Theorem 1.3 ([7]).
Let $\mathcal{H}$ be a computable projection-closed class of graphs. Then the following holds for edge-weighted $\mathcal{H}$).
  1. 1.
    If every graph in $\mathcal{H}$ is a trivial pattern, then there is a polynomial-time algorithm.
  2. 2.
    Otherwise, if $\mathcal{H}$ has bounded distance to extended bicliques, then
    1. (a)
      There is an $f(\pi,t)n^{O(\sqrt{\pi})}$ time algorithm.
    2. (b)
      Assuming ETH, there is a universal constant $\alpha>0$ such that for any fixed choice of $\pi\geq 0$, there is no $O(n^{\alpha\sqrt{\pi}})$ algorithm, even when restricted to instances in $\textup{\sf planar}+\pi e$ with $t=3$ terminals.
  3. 3.
    Otherwise,
    1. (a)
      There is an $f(\pi,t)n^{O(\sqrt{\pi+t})}$ algorithm.
    2. (b)
      Assuming ETH, there is a universal constant $\alpha>0$ such that for any fixed choice of $\pi\geq 0$ and $t\geq 3$, there is no $O(n^{\alpha\sqrt{\pi+t}})$ algorithm, even when restricted to $\textup{\sf planar}+\pi e$ instances with at most $t$ terminals.
Observe that $\pi$ has a different form of influence on the exponent compared to genus: it is $O(\sqrt{\pi})$ instead of $O(g)$ in the case of bounded distance to extended bicliques, and $O(\sqrt{\pi+t})$ instead of $O(\sqrt{g^{2}+gt+t})$ in the unbounded distance case. Intuitively, having $\pi$ extra edges is a much more restricted extension of planar graphs than having genus $g$, and consequently, the parameter $\pi$ has a much weaker influence on the exponent than genus has.
For unweighted graphs, Focke et al. [7] showed that the problem is easier than the edge-weighted version and the parameter $\pi$ can be removed from the exponent.
Theorem 1.4 ([7]).
Unweighted Multicut can be solved in time $f(\pi,t)n^{O(\sqrt{t})}$.
However, this result leaves open the question what happens in unweighted $\mathcal{H}$) if $\mathcal{H}$ has bounded distance to extended bicliques. Based on Theorems 1.2 and 1.3, one would expect that it is possible to remove the number $t$ of terminals from the exponent as well.
Our contributions
First, we resolve the remaining question about unweighted $\mathcal{H}$) and show that if $\mathcal{H}$ has bounded distance to extended bicliques, then both $t$ and $\pi$ can be removed from the exponent. Recall that $\mu$ is the number of vertices needed to be deleted to make the demand graph an extended biclique. Our first main result is the following algorithm.
Theorem 1.5.
Unweighted Multicut can be solved in time $f(\pi,t)n^{O(\sqrt{\mu})}$.
Observe that, in contrast to Theorem 1.4, only $\mu$ is in the exponent of the running time of the algorithm of Theorem 1.5. As $\mu\leq t$, this algorithm gives an independent proof of Theorem 1.4 with a very different proof technique. With Theorem 1.5 at hand, we can complete the picture for the complexity of unweighted $\mathcal{H}$):
Theorem 1.6.
Let $\mathcal{H}$ be a computable projection-closed class of graphs. Then the following holds for unweighted $\mathcal{H}$).
  1. 1.
    If $\mathcal{H}$ has bounded distance to extended bicliques, then there is an $f(\pi,t)n^{O(1)}$-time algorithm.
  2. 2.
    Otherwise,
    1. (a)
      There is an $f(\pi,t)n^{O(\sqrt{t})}$ algorithm.
    2. (b)
      Assuming ETH, there is a universal constant $\alpha>0$ such that for any fixed choice of $t\geq 3$, there is no $O(n^{\alpha\sqrt{t}})$-time algorithm, even when restricted to planar instances with at most $t$ terminals.
We consider yet another, even more restricted parameter measuring the distance from planar graphs: the crossing number, the minimum number of crossings needed when drawing the graph in the plane. Clearly, if a graph has a drawing with $cr$ crossings, then we can remove a set of at most $cr$ edges to make the graph planar. Thus the crossing number $cr$ is always at least $\pi$, that is, an algorithm parameterized by $\pi$ immediately gives an algorithm parameterized by $cr$.
Again, there is no straightforward reduction from the edge-weighted version to the unweighted version without increasing the crossing number. Thus both versions of the problem when parameterizing by crossing number are interesting. Clearly, the unweighted version is no harder than the edge-weighted one. Moreover, both versions parameterized by crossing number are no harder than their respective versions parameterized by $\pi$. But which of these versions are strictly harder than the others?
It turns out that, when parameterizing by crossing number, the edge weights do not change the hardness of the problem and both the edge-weighted and the unweighted versions parameterized by $cr$ have the same complexity as the unweighted version parameterized by $\pi$. The following algorithm is our main result on the crossing number parameterization.
Theorem 1.7.
Multicut can be solved in time $f(cr,t)n^{O(\sqrt{\mu})}$.
Together with Theorem 1.1, it implies the following complete classification for the complexity of $\mathcal{H}$) when parameterizing by $cr$:
Theorem 1.8.
Let $\mathcal{H}$ be a computable projection-closed class of graphs. Then the following holds for edge-weighted $\mathcal{H}$).
  1. 1.
    If $\mathcal{H}$ has bounded distance to extended bicliques, then there is an $f(cr,t)n^{O(1)}$-time algorithm.
  2. 2.
    Otherwise,
    1. (a)
      There is an $f(cr,t)n^{O(\sqrt{t})}$ algorithm.
    2. (b)
      Assuming ETH, there is a universal constant $\alpha>0$ such that for any fixed choice of $t\geq 3$, there is no $O(n^{\alpha\sqrt{t}})$-time algorithm, even when restricted to unweighted planar instances with at most $t$ terminals.
Note that the algorithmic results of Theorems 1.6 and 1.8 are incomparable: Theorem 1.6 considers only unweighted graphs, while Theorem 1.8 considers edge-weighted graphs, but under parameterization by a potentially much larger parameter.
Theorems 1.21.8 complete the picture of how different parameters measuring closeness to planarity influence the running time of $\mathcal{H}$) for different classes $\mathcal{H}$ (see Figure 1). Let us observe that the exact form of the influence is very different for the different parameters. As we explain below, the treewidth of the so-called multicut dual of the optimum solution is the main factor determining the exponent of the running time [3, 7]. The exponent $O(\sqrt{t})$ for planar graphs comes from the fact that a planar graph with $O(t)$ vertices (or faces) has treewidth $O(\sqrt{t})$. Intuitively, for nonplanar graphs, the situation changes as follows:
  • Genus. If the genus is large, then a graph with $t$ vertices can have treewidth larger than $O(\sqrt{t})$. The maximum possible treewidth is given by the (nonobvious) formula $O(\sqrt{g^{2}+gt+t})$. Therefore, the maximum treewidth of the multicut dual is $O(\sqrt{g^{2}+gt+t})$, giving a bound on the exponent of the running time.
  • Edge-weighted $\textup{\sf planar}+\pi e$. With extra edges having infinite (or very large) weight, we can simulate multiple terminals at different parts of a planar graph. More precisely, a planar instance of Group 3-Terminal Cut with $t$ terminals in each of the three groups can be reduced to an instance of Multiway Cut with three terminals if we connect the terminals in each group with a total of $\pi=3(t-1)$ infinite-weight edges, creating a $\textup{\sf planar}+\pi e$ instance. Thus these extra edges influence the running time exactly as extra terminals would do.
  • Unweighted $\textup{\sf planar}+\pi e$. We can always try to solve the problem by putting the $\pi$ extra edges into the solution, removing them from $G$, and then solving the resulting planar instance. Of course, this might not always lead to an optimum solution, but can be up to $\pi$ edges larger. However, this argument allows us to say that if the solution contains a cut that is “inefficient” in the sense that it is larger then the minimum cut by more than $\pi$, then the cut can be replaced by a minimum cut and the $\pi$ extra edges, resulting in a srictly smaller solution. Using these type of arguments, the problem can be reduced to a planar setting where the number of faces of the multicut dual does not depend on $\pi$.
  • Crossing number. Each crossing can increase the number of vertices of the multicut dual by one, thus the number of vertices and the treewidth of the multicut dual can be $O(t+cr)$ and $O(\sqrt{t+cr})$, respectively. However, all these extra vertices of the multicut dual are at fixed places determined by the crossing. As pointed out by Focke et al. [8] it is sufficient to consider the treewidth of the multicut dual after removing all the vertices whose locations are fixed (see Theorem 2.3). Indeed, after removing these vertices, the treewidth of the multicut dual does not depend on the crossing number.
Some remarks on two further notion of almost-planar graphs can be found in Appendix B.
weighted unweighted crossing
genus $\textup{\sf planar}+\pi e$ $\textup{\sf planar}+\pi e$ number
$\mathcal{H}$ with
bounded distance $f(g,t)n^{{\color[rgb]{1,0,0}O(g)}}$ $f(\pi,t)n^{{\color[rgb]{1,0,0}O(\sqrt{\pi})}}$ $f(\pi,t)n^{{\color[rgb]{1,0,0}O(1)}}$ (*) $f(cr,t)n^{{\color[rgb]{1,0,0}O(1)}}$ (*)
to extended bicliques
$\mathcal{H}$ with
unbounded distance $f(g,t)n^{{\color[rgb]{1,0,0}O(\sqrt{g^{2}+gt+t})}}$ $f(\pi,t)n^{{\color[rgb]{1,0,0}O(\sqrt{\pi+t})}}$ $f(\pi,t)n^{{\color[rgb]{1,0,0}O(\sqrt{t})}}$ $f(cr,t)n^{{\color[rgb]{1,0,0}O(\sqrt{t})}}$ (*)
to extended bicliques
Figure 1: The running time of the algorithms for $\mathcal{H}$) with the different parameterizations. The new results of the paper are marked by (*). Note that, for every fixed projection-closed class $\mathcal{H}$, the exponents of the running time are tight (up to logarithmic factors).
2 Preliminaries
Some basic notation on sets, graphs, and multicuts can be found in Appendix A. We now give some less well-known, problem-specific definitions.
An extended biclique decomposition of a graph $H$ is a partition $(B_{1},B_{2},I,X)$ of $V(H)$ such that all vertices of $I$ are isolated in $H-X$ and $E(H)$ contains the edge $t_{1}t_{2}$ for all $t_{1}\in B_{1}$ and $t_{2}\in B_{2}$. We use $\mathcal{H}_{\mu}$ for the set of graphs $H$ that admit an extended biclique decomposition $(B_{1},B_{2},I,X)$ with $|X|\leq\mu$.
We now review the concept of so-called multicut duals, which will play a crucial role throughout the article. Recall that an instance $(G,H)$ of Multicut is planar if $G$ is given in the form of a graph cellularly embedded in the plane. Given a graph $C$ that is embedded in the plane and in general position with $G$, we denote by $e_{G}(C)$ those edges of $G$ that are crossed by at least one edge of $C$. The weight $w(C)$ of $C$ is defined to be $w(e_{G}(C))$ where $w$ is the weight function associated with $G$. A multicut dual for $(G,H)$ is a graph $C$ embedded on $N$ that is in general position with $G$ and has the property that, for every $t_{1},t_{2}\in V(H)$ with $t_{1}t_{2}\in E(H)$, the terminals $t_{1}$ and $t_{2}$ are contained in different faces of $C$.
The following result is an immediate consequence of Lemmas 2.3 and 2.4 in [8].
Lemma 2.1.
Let $(G,H)$ be a planar instance of Multicut and $C$ a minimum weight multicut dual for $(G,H)$. Then $e_{G}(C)$ is a minimum weight multicut for $(G,H)$.
In order to make use of the topology of certain multicut duals, we need two results related to treewidth. The first one of the following two results provides a sufficient criterion for the treewidth of certain graphs being bounded and the second one allows to exploit bounded treewidth algorithmically.
The following result can be obtained from Theorem 3.2 in [6] and the fact that the treewidth of any graph can only be larger than its branchwidth by a constant factor.
Proposition 2.2.
Let $C$ be a planar graph that admits a $5$-dominating set of size $k$ for some positive integer $k$. Then $\textup{tw}(C)=O(\sqrt{k})$.
An extended planar instance $(G,H,F^{*})$ of Multicut is a planar instance $(G,H)$ of Multicut together with a set $F^{*}$ of faces of $G$ which is given with the input. Given a set $F^{*}$ of faces of $G$ and a multicut dual $C$ for $(G,H)$, let $C-F^{*}$ be the graph obtained from $C$ by removing every vertex that is in a face of $F^{*}$ and removing every edge that intersects a face of $F^{*}$. The following result is the restriction of Theorem 1.11 in [8] to the planar setting.
Theorem 2.3.
Let $(G,H,F^{*})$ be an extended planar instance of Multicut such that every subcubic inclusionwise minimal minimum-weight multicut dual $C$ satisfies $\textup{tw}(C-F^{*})\leq\beta$. Then an optimum solution of $(G,H)$ can be found in time $f(|F^{*}|,t)n^{O(\beta)}$.
3 Edge-deletion distance to planarity
In this section, we give an overview of the proof of the following result. See 1.5 Throughout this section, we say that an instance $(G,H)$ of Multicut is $\textup{\sf planar}+\pi e$ if a set $E_{\pi}\subseteq E(G)$ is given with the instance such that $|E_{\pi}|=\pi$ and $G\setminus E_{\pi}$ is planar. We further use $T$ for $V(H)$, $W$ for $V(E_{\pi})$, and $G_{0}$ for $G\setminus E_{\pi}$. Also, for some $S\subseteq E(G)$, we use $S_{0}$ for $S\setminus E_{\pi}$. We say that $(G,H)$ is connected if $G$ is connected.
The overall algorithmic strategy for Theorem 1.5 is to run an algorithm on the given instance that either produces a minimum multicut for the instance or computes a collection of instances for which the parameters in consideration are smaller and such that from minimum multicuts of all those instances, a minimum multicut for the original instance can be computed. We need the following definitions.
Given instances $(G,H)$ and $(G^{\prime},H^{\prime})$ of Multicut, $(G^{\prime},H^{\prime})$ is a subinstance of $(G,H)$ if $G^{\prime}$ and $H^{\prime}$ are labelled subgraphs of $G$ and $H$, respectively. Next, an extended subinstance of $(G,H)$ consists of a set $S^{\prime}\subseteq E(G)$ and a subinstance $(G^{\prime},H^{\prime})$ of $(G,H)$ with $E(G^{\prime})\cap S^{\prime}=\emptyset$. When we say that an extended subinstance $(S^{\prime},(G^{\prime},H^{\prime}))$ has certain properties which can only be associated to an instance of Multicut, we refer to $(G^{\prime},H^{\prime})$. Next, we say that an extended subinstance $(S^{\prime},(G^{\prime},H^{\prime}))$ of $(G,H)$ is optimumbound for $(G,H)$ if $S^{\prime}\cup S^{\prime\prime}$ is a minimum multicut for $(G,H)$ for every minimum multicut $S^{\prime\prime}$ for $(G^{\prime},H^{\prime})$.
The main technical difficulty is to prove the following result, which shows that the above described reduction to smaller instances exists for connected instances. With this result at hand, Theorem 1.5 follows readily.
Lemma 3.1.
Let $(G,H)$ be a connected $\textup{\sf planar}+\pi e$ instance of Multicut. Then, in $f(\pi,t)n^{O(\sqrt{\mu})}$, we can run an algorithm that returns a set $S\subseteq E(G)$ and a collection $(S_{i},(G_{i},H_{i}))_{i\in[k]}$ of extended subinstances of $(G,H)$ for some $k=f(\pi,t)$ such that for all $i\in[k]$, we have that $\pi(G_{i})+|V(H_{i})|<\pi(G)+|V(H)|$ holds and either $S$ is a minimum multicut for $(G,H)$ or there exists some $i\in[k]$ such that $(S_{i},(G_{i},H_{i}))$ is optimumbound for $(G,H)$.
We now give an overview of the proof of Lemma 3.1. Let $(G,H)$ be a $\textup{\sf planar}+\pi e$ instance of Multicut. The idea is to have a branching algorithm that gradually guesses more and more crucial information about a minimum multicut for $(G^{\prime},H^{\prime})$ while the number of directions it branches into remains bounded. The current status of this procedure is displayed by a subpartition of the vertices of $V(G)$ with some extra properties. More formally, a state is a subpartition $\mathcal{P}$ of $V(G)$ with $W\cup T\subseteq\bigcup\mathcal{P}$ and $Y\cap(W\cup T)\neq\emptyset$ for all $Y\in\mathcal{P}$. Given a certain state, the objective is to understand whether there exists a minimum multicut $S$ for $(G,H)$ such that the structure of the components of $G_{0}\setminus S$ reflects the structure of $\mathcal{P}$. In order to make this more formal, we need the following definitions: Given a set $V$ and two subpartitions $\mathcal{P}^{1}$ and $\mathcal{P}^{2}$ of $V$, we say that $\mathcal{P}^{2}$ extends $\mathcal{P}^{1}$ if $|\mathcal{P}^{1}|=|\mathcal{P}^{2}|$ and for every $Y\in\mathcal{P}^{1}$, there exists some $\overline{Y}\in\mathcal{P}^{2}$ with $Y\subseteq\overline{Y}$ such that for any distinct $Y_{1},Y_{2}\in\mathcal{P}^{1}$, we have $\overline{Y_{1}}\neq\overline{Y_{2}}$. Given a set $S\subseteq E(G_{0})$, we use $\mathcal{K}(G_{0}\setminus S_{0})$ for the unique partition of $V(G)$ in which two vertices are in the same partition class if and only if they are in the same component of $G_{0}\setminus S_{0}$. For a state $\mathcal{P}$, we say that a multicut $S$ for $(G,H)$ respects $\mathcal{P}$ if $\mathcal{K}(G_{0}\setminus S_{0})$ extends $\mathcal{P}$ and we say that a state is valid if there exists a minimum multicut for $(G,H)$ respecting it. Further, $\mathcal{P}$ is maximum valid if $|\mathcal{P}|$ is maximum among all valid states.
The idea of our algorithm now consists of considering a valid state, choosing a vertex that is not contained in any class of the state and making the classes larger by adding this vertex to one of them. As we do not know which class this vertex should be added to, we try all classes of the state and branch in these directions. The difficulty is now to carefully choose the vertex in consideration and make some additional modifications on the states so that this procedure terminates sufficiently fast.
In order to be able to explain how this works, we first need some more definitions. Given disjoint sets $Y_{1},Y_{2}\subseteq V(G)$, we use $\lambda_{G_{0}}(Y_{1},Y_{2})$ to denote the size of the smallest set of edges in $E(G_{0})$ whose deletion from $G_{0}$ results in a graph in which there exists no component containing a vertex of $Y_{1}$ and a vertex of $Y_{2}$. Next, for a state $\mathcal{P}$ and some $Y\in\mathcal{P}$, we define $\alpha_{G,\mathcal{P}}(Y)=\lambda_{G_{0}}(Y,\bigcup\mathcal{P}\setminus Y)-\lambda_{G_{0}}(Y\cap(W\cup T),W\cup T\setminus Y)$. Intuitively speaking, this measure describes how much more costly it is to separate the class from the other classes of the state in comparison to only separating the corresponding elements of $W\cup T$. Given a state $\mathcal{P}$, we now classify its classes according to their value of $\alpha$ and the value of $\alpha$ of their neighboring classes. More formally, given a state $\mathcal{P}$ and some $Y\in\mathcal{P}$, we say that $Y$ is fat in $\mathcal{P}$ if $\alpha_{G_{0},\mathcal{P}}(Y)\geq\pi(t+1)+1$. For some $Y\in\mathcal{P}$ that is not fat, we say that $Y$ is fat-neighboring in $\mathcal{P}$ if there exists some fat $Y^{\prime}\in\mathcal{P}$, some $u\in Y$, and some $v\in Y^{\prime}$ such that $uv\in E(G_{0}).$ Finally, for some $Y\in\mathcal{P}$ which is neither fat nor fat-neighboring in $\mathcal{P}$, we say that $Y$ is thin in $\mathcal{P}$. We refer by $\mathcal{P}_{\textup{fat}},\mathcal{P}_{\textup{f-n}}$, and $\mathcal{P}_{\textup{thin}}$ to the fat, fat-neighboring, and thin classes of $\mathcal{P}$, respectively.
In order to make use of this distinction, we crucially need a structural property of minimum multicuts for $(G,H)$. Namely, the following result shows that, for any minimum multicut $S$ for $(G,H)$ and any vertex set of a component of $G_{0}\setminus S_{0}$, one of two things holds: either the value of $\alpha$ is not too large or a terminal of $X$ is close to this partition class in a certain sense, where $(B_{1},B_{2},I,X)$ is an extended biclique decomposition of $H$. In order to measure this proximity of certain vertex sets with respect to a multicut, we introduce the following notion: Let $S\subseteq E(G)$, and $Y_{1},Y_{2}\subseteq V(G)$. Then we denote by $\textup{dist}_{G_{0},S_{0}}(Y_{1},Y_{2})$ the smallest integer $q$ such that $G_{0}$ contains a $Y_{1}Y_{2}$-path $P$ with $|E(P)\cap S_{0}|=q$. We are now ready to state the following structural result for minimum multicuts.
Lemma 3.2.
Let $(G,H)$ be a $\textup{\sf planar}+\pi e$ instance of Multicut, let $S$ be a minimum multicut for $(G,H)$ and let $Y\in\mathcal{P}_{\textup{fat}}$ for $\mathcal{P}=\mathcal{K}(G_{0}\setminus S_{0})$. Then $\textup{dist}_{G_{0},S_{0}}(Y,X)\leq 1$.
We now explain the strategy of our algorithm. Given a state with an extra property that will be explained later, we choose a vertex that is not contained in any class of the state and has a neighbor in a thin class of the state in $G_{0}$. We then guess a partition class which this vertex will be added to and modify the state accordingly. In order to show the functionality of this algorithm, there are two points that need further explanation. First, we need to explain how to conclude when the described modification is no longer possible and second, we need to explain why this procedure terminates sufficiently fast.
We first explain the scenario that we can no longer execute the described modification, so there does not exist any vertex in the graph that is not contained in any class of the state $\mathcal{P}$ in consideration and that has a neighbor in a thin class of $\mathcal{P}$ in $G_{0}$. We call such a state complete, all other states are called incomplete. For complete states, we need to consider two different cases, depending on whether $\mathcal{P}$ contains a thin class.
The case that there exists a thin class $Y$ can be handled with elementary means. Observe that for this thin class, all its neighbors in $G_{0}$ are contained in a different class of $\mathcal{P}$. Hence, if $\mathcal{P}$ is valid, then all edges in $\delta_{G_{0}}(Y)$ are contained in a minimum multicut for $(G,H)$. We can hence delete these edges and solve the remaining instance, which turns out to be significatly smaller in a certain sense. This is subsumed in the following result.
Lemma 3.3.
Let $(G,H)$ be a connected $\textup{\sf planar}+\pi e$ instance of Multicut and let a complete valid state $\mathcal{P}$ of $(G,H)$ with $\mathcal{P}_{\textup{thin}}\neq\emptyset$ be given. Then in polynomial time, we can compute an optimumbound extended subinstance $(S^{\prime},(G^{\prime},H^{\prime}))$ of $(G,H)$ such that $\pi(G^{\prime})+|V(H^{\prime})|<\pi(G)+|V(H)|$.
We now turn to the case that $\mathcal{P}$ does not contain a thin set. In order to handle this case, geoemetrical arguments on multicut duals will be needed. Namely, we show that if $\mathcal{P}$ is valid and does not contain a thin set, then the problem of finding a minimum multicut for $(G,H)$ can be reduced to solving a planar instance of multicut, which can be solved sufficiently fast due to Theorem 2.3 and an argument on the treewidth of the multicut duals for this instance. Formally, we prove the following result, where $\tau(\mathcal{P})$ denotes the sum of the number of components of $G_{0}[Y]$ over all $Y\in\mathcal{P}$.
Lemma 3.4.
Let $(G,H)$ be a connected $\textup{\sf planar}+\pi e$ instance of Multicut and let a complete maximum valid state $\mathcal{P}$ of $(G,H)$ with $\mathcal{P}_{\textup{thin}}=\emptyset$ be given. Then a minimum multicut for $(G,H)$ can be computed in $f(\pi,\tau(\mathcal{P}))n^{O(\sqrt{\mu})}$.
It remains to show that we can design our procedure so that we reach a complete state significantly fast. In order to keep track of the progress of our algorithm, we need a measure of how far advanced a given state is. To this end, given a state $\mathcal{P}$, we first define an integer $\kappa(\mathcal{P},Y)$ for every $Y\in\mathcal{P}$ by
$\kappa(\mathcal{P},Y)=\left\{\begin{array}[]{lr}2(\pi(t+1)+1),&\text{for }Y\in\mathcal{P}_{\textup{fat}},\\ \pi(t+1)+1,&\text{for }Y\in\mathcal{P}_{\textup{f-n}},\\ \alpha_{G_{0},\mathcal{P}}(Y),&\text{for }Y\in\mathcal{P}_{\textup{thin}}.\end{array}\right\}$
Now, we define $\kappa(\mathcal{P})=\sum_{Y\in\mathcal{P}}\kappa(\mathcal{P},Y)$.
Our objective is to make the value of $\kappa$ increase in each iteration of the algorithm. In order to make this possible, we need that the addition of a vertex to a given set in the state makes the value of $\alpha$ of this set increase. In order for that to hold, we need the classes in the state to be as large as possible in a certain sense. More precisely, given a graph $G$ and two disjoint sets $Y_{1},Y_{2}\subseteq V(G)$, we say that a set $Y_{3}$ is relevant for $(Y_{1},Y_{2})$ in $G$ if $Y_{1}\subseteq Y_{3},Y_{2}\cap Y_{3}=\emptyset,d_{G}(Y_{3})=\lambda_{G}(Y_{1},Y_{2})$ and $Y_{3}$ is maximum among all sets with these properties. Further, we say that a state $\mathcal{P}$ is relevant if for all $Y\in\mathcal{P}$, we have that $Y$ is a relevant set for $(Y,\bigcup\mathcal{P}\setminus Y)$. Luckily, the following result whose proof needs some careful reconfiguration arguments, allows us to turn arbitrary states in consideration into relevant ones without dropping any important properties.
Lemma 3.5.
Let a connected $\textup{\sf planar}+\pi e$ instance $(G,H)$ of Multicut and a state $\mathcal{P}$ of $(G,H)$ be given. Then, in polynomial time, we can compute a relevant state $\mathcal{P}^{\prime}$ of $(G,H)$ such that $|\mathcal{P}^{\prime}|=|\mathcal{P}|,\tau(\mathcal{P}^{\prime})\leq\tau(\mathcal{P})$, $\kappa(\mathcal{P}^{\prime})\geq\kappa(\mathcal{P})$ and, if $\mathcal{P}$ is maximum valid, then so is $\mathcal{P}^{\prime}$.
With this result at hand, we can handle incomplete states. As pointed out before, given an incomplete relevant state $\mathcal{P}$, we choose a vertex $u_{0}\in V(G)\setminus\bigcup\mathcal{P}$ that has a neighbor in a thin class $Y$ of $\mathcal{P}$. We now consider all states that are obtained by adding $u_{0}$ to one of these partition classes. If $u_{0}$ is added to a thin or a fat-neighboring partition class, then, as $\mathcal{P}$ is relevant, the value of $\alpha$ increases for this class and hence $\kappa$ increases. If $u_{0}$ is added to a fat class, then $Y$ becomes fat-neighboring and hence $\kappa$ also increases. This shows that a complete state can be reached sufficiently fast. In order to apply Lemma 3.4, we also need to make sure that $\tau(\mathcal{P})$ does not increase too fast. However, this follows readily by construction. Our handling of relevant incomplete states is subsumed in the following result.
Lemma 3.6.
Let $(G,H)$ be a connected $\textup{\sf planar}+\pi e$ instance of Multicut and let a relevant incomplete state $\mathcal{P}$ of $(G,H)$ be given. Then, in polynomial time, we can compute a collection $(\mathcal{P}^{i})_{i\in[q]}$ of $q$ states of $(G,H)$ for some $q\leq\pi+t$ such that $\tau(\mathcal{P}^{i})\leq\tau(\mathcal{P})+1$ and $\kappa(\mathcal{P}^{i})>\kappa(\mathcal{P})$ hold for all $i\in[q]$ and if $\mathcal{P}$ is maximum valid, then there exists some $i\in[q]$ such that $\mathcal{P}^{i}$ is maximum valid.
The overall strategy for handling Lemma 3.1 is now to apply Lemma 3.5 and Lemma 3.6 until we are left with a collection of complete states. These can then be handled by Lemma 3.3 and Lemma 3.4.
4 Crossing number
In this section, an embedding of a graph $G$ consists of a mapping of the vertices of $G$ to points in the plane and a mapping of the edges of $G$ to curves in the plane connecting their endpoints such that all vertices and edges of $G$ are in general position except the required adjacencies and such that in any point not corresponding to a vertex, at most two edges intersect. An edge crossing refers to a point that does not correspond to a vertex and in which two edges intersect. Given an instance $(G,H)$ of Multicut  we use $cr$ for the crossing number of a graph $G$, which is defined to be the smallest integer $k$ such that there exists an embedding of $G$ with exactly $k$ edge crossings. The following is the main result of this section. See 1.7 In order to make our algorithm work, rather than just the crossing number of a graph, we also need to have an optimal embedding available. This can be achieved through the following result due to Grohe [9].
Proposition 4.1.
Given a graph $G$, there exists an algorithm that computes an embedding of $G$ in the plane minimizing the number of crossings and that runs in time $f(cr)n^{O(1)}$.
Formally, an instance $(G,H)$ of Multicut is called crossing-planar if $G$ is given as an embedding of a graph. Given a crossing-planar instance, we denote by $E_{cr}$ the set of edges in $E(G)$ crossing at least one other edge of $E(G)$ and we set $\overline{cr}=|E_{cr}|$.Observe that $\overline{cr}$ can be arbitrarily large in comparison to $cr$ as $\overline{cr}$ refers to the given embedding and $cr$ to an optimal embedding of the abstract graph $G$. However, when considering an optimal embedding of $G$, we have $\overline{cr}\leq 2cr$ as every crossing involves 2 edges.
For our algorithm, it will be convenient to assume that the instances in consideration satisfy a technical extra condition. Namely, a crossing-planar instance $(G,H)$ of Multicut is called normalized if every $e\in E_{cr}$ is of infinite weight. Observe that for any normalized crossing-planar instance $(G,H)$ of Multicut, any finite multicut $S$ of $(G,H)$ and any $e\in E_{cr}$, we have that $V(e)$ is contained in one component of $G\setminus S$. Our main technical contribution is the proving the restriction of Theorem 1.7 to normalized instances.
Theorem 4.2.
There exists an algorithm that solves every normalized crossing-planar instance of Multicut and runs in $f(\overline{cr},t)n^{O(\sqrt{\mu})}$.
The simple proof that Theorem 4.2 implies Theorem 1.7 is postponed to Appendix D. It remains to prove Theorem 4.2.
Given a normalized crossing-planar instance $(G,H)$ of Multicut, we denote by $\mathcal{H}$ the set of graphs $H^{\prime}$ on $V(H)\cup V(E_{cr})$ and we use $G^{\prime}$ for $G\setminus E_{cr}$. Further, we let $F^{*}$ denote the faces of $G^{\prime}$ which contain a pair of crossing edges in $G$. The main technical difficulty for the proof of Theorem 4.2 is contained in the following result.
Lemma 4.3.
Let $(G,H)$ be a normalized crossing-planar instance of Multicut admitting a multicut of finite weight. Then there exists $H^{\prime}\in\mathcal{H}$ such that every minimum weight multicut for $(G^{\prime},H^{\prime})$ is a minimum weight multicut for $(G,H)$ and for every subcubic, inclusion-wise minimal minimum weight multicut dual $C$ for $(G^{\prime},H^{\prime})$, we have $\textup{tw}(C-F^{*})=O(\sqrt{\mu})$.
Due to space restrictions, the full proof of Lemma 4.3 is postponed to Appendix F. We now give an overview of the proof of Lemma 4.3.
Let $(B_{1},B_{2},I,X)$ be an extended biclique decomposition of $H$ with $|X|=\mu$. Further, let $S^{*}$ be a minimum weight multicut for $(G,H)$ such that $S^{*}$ minimizes the number of elements of $V(H)\cup V(E_{cr})$ that are in a component of $G\setminus S^{*}$ that also contains a terminal of $X$. In the following, we use $\overline{X}$ for the set of elements of $V(H)\cup V(E_{cr})$ that are in a component of $G\setminus S^{*}$ that also contains a terminal of $X$, we use $\overline{B_{1}}$ for the set of elements of $(V(H)\cup V(E_{cr}))\setminus\overline{X}$ that are in a component of $G\setminus S^{*}$ that also contains a terminal of $B_{1}$, and we use $\overline{B_{2}}$ for $(V(H)\cup V(E_{cr}))\setminus(\overline{X}\cup\overline{B_{1}})$. Observe that $(\overline{B_{1}},\overline{B_{2}},\overline{X})$ is a partition of $V(H)\cup V(E_{cr})$. We now define $H^{\prime}$ to be the complete multipartite graph such that $\overline{B_{1}}$ and $\overline{B_{2}}$ are partition classes and such that for every component of $G\setminus S^{*}$ that contains at least one terminal of $X$, all the elements of $\overline{X}$ contained in this component form a partition class.
This finishes the description of $H^{\prime}$. Observe that $H^{\prime}\in\mathcal{H}$. Further observe that $H$ is a subgraph of $H^{\prime}$.The following two claims directly imply that every minimum weight multicut for $(G^{\prime},H^{\prime})$ is also a minimum weight multicut for $(G,H^{\prime})$. Claim 4.4 will be reused for the proof of the treewidth bound. These claims mainly follow from the construction of $(G^{\prime},H^{\prime})$.
Claim 4.4.
Let $S^{\prime}$ be a multicut for $(G^{\prime},H^{\prime})$. Then $S^{\prime}$ is also multicut for $(G,H^{\prime})$.
Claim 4.5.
$S^{*}$ is a multicut for $(G^{\prime},H^{\prime})$.
We are now ready to conclude that any minimum weight multicut $S$ for $(G^{\prime},H^{\prime})$ is a minimum weight multicut for $(G,H)$. First observe that $S$ is a multicut for $(G,H)$ by Claim 4.4 and as $H$ is a subgraph of $H^{\prime}$. Further, by Claim 4.5 and the minimality of $S$, we obtain $w(S)\leq w(S^{*})$ where $w$ is the weight function associated to $G$. By the minimality of $S^{*}$, the statement follows.
In the following, let $C$ be a subcubic, inclusion-wise minimal, minimum weight multicut dual for $(G^{\prime},H^{\prime})$. We will give the desired treewidth bound on $C$. Let $S^{\prime}=E_{G}(C)$ and observe that $S^{\prime}$ is a minimum weight multicut for $(G^{\prime},H^{\prime})$ by Lemma 2.1. It hence follows from the first part of the lemma that $S^{\prime}$ is also a minimum weight multicut for $(G,H)$. Further, we use $C^{\prime}$ for $C-F^{*}$.
Our strategy for proving the second part of the lemma is to show that the faces of $C^{\prime}$ containing a terminal of $\overline{X}$ form a small set of faces that is close to every face of $C^{\prime}$ in a certain sense. We first give the following result showing that the number of such faces is small indeed. Its proof uses the fact that $S^{*}$ is chosen so that the size of $\overline{X}$ is minimized.
Claim 4.6.
Let $f$ be a face of $C^{\prime}$ that contains a terminal of $\overline{X}$. Then $f$ also contains a terminal of $X$.
The next result, uses the structure of $H$.
Claim 4.7.
Let $v\in V(C^{\prime})$ with $d_{C^{\prime}}(v)=3$. Then $v$ is incident to a face of $C^{\prime}$ that contains a terminal of $\overline{X}$.
We are now ready to conclude the desired treewidth bound on $C^{\prime}$. Let $C^{\prime\prime}$ be a component of $C^{\prime}$. For every face $f$ of $C^{\prime\prime}$, we now introduce a vertex $z_{f}$ that shares an edge with every vertex that is incident to $f$. Let $C^{\prime\prime}_{0}$ be the resulting graph. It turns out that $C^{\prime\prime}_{0}$ is a planar graph and the vertices corresponding to faces containing terminals of $\overline{X}$ form a 3-dominating set in $C^{\prime\prime}_{0}$. By Proposition 2.2 and $|A|\leq|X|\leq\mu$, we obtain that $\textup{tw}(C^{\prime\prime}_{0})=O(\sqrt{\mu})$ and hence $\textup{tw}(C^{\prime})=O(\sqrt{\mu})$. This allows to conclude Lemma 4.3.
With Lemma 4.3 at hand, it is easy to conclude Theorem 4.2. The proof can be found in Appendix E.
Appendix A Basic notation
We here give some basic notation on sets, graphs, and multicuts.
For some positive integer $k$, we use $[k]$ for $\{1,\ldots,k\}$. A subpartition of a set $V$ is a collection $\mathcal{P}$ of disjoint subsets of $V$. We use $\bigcup\mathcal{P}$ for the union of all classes of $\mathcal{P}$, Moreover, $\mathcal{P}$ is a partition if $\bigcup\mathcal{P}=V$.
A graph $G$ consists of a finite set $V(G)$ whose elements are called vertices and a set $E(G)$ of subsets of $V(G)$ of size 2 called edges. For some $E^{\prime}\subseteq E(G)$, we use $V(E^{\prime})$ for the set of vertices contained in at least one edge in $E^{\prime}$. Further, we use $G\setminus E^{\prime}$ for the unique graph on $V(G)$ with $E(G\setminus E^{\prime})=E(G)\setminus E^{\prime}$. Next, for some $V^{\prime}\subseteq V(G)$, we use $G[V^{\prime}]$ for the subgraph of $G$ induced by $V^{\prime}$ that is, $V(G[V^{\prime}])=V^{\prime}$ and $E(G[V^{\prime}])$ contains all edges of $E(G)$ fully contained in $V^{\prime}$. We also use $G-V^{\prime}$ for $G[V(G)\setminus V^{\prime}]$ Given a graph $G$ and two disjoint sets $X,Y\subseteq V(G)$, we use $\delta_{G}(X,Y)$ for the set of edges in $E(G)$ containing a vertex of $X$ and a vertex of $Y$. We further use $d_{G}(X,Y)$ for $|\delta_{G}(X,Y)|,\delta_{G}(X)$ for $\delta_{G}(X,V(G)\setminus X)$, and $d_{G}(X)$ for $|\delta_{G}(X)|$. We say that $G$ is connected if $d_{G}(X)\geq 1$ for every $X\subseteq V(G)$. Next, we say that a graph $G^{\prime}$ is a subgraph of $G$ if $V(G^{\prime})\subseteq V(G)$ and $E(G^{\prime})\subseteq E(G)$. A component of $G$ is a maximal connected subgraph of $G$.
A graph $P$ is a $u_{1}u_{2}$-path for some $u_{1},u_{2}\in V(P)$ if $P$ is connected, $d_{P}(u_{i})=1$ holds for all $i\in[2]$ and $d_{P}(v)=2$ holds for all $v\in V(P)$. We say that the length of $P$ is $|E(P)|$. If $u_{2}\in U$ for some $U\subseteq V(G)\setminus\{u_{1}\}$, we say that $P$ links $u_{1}$ and $U$. For the first of these results, for a graph $G$ and a positive integer $k$, we say that a set $X\subseteq V(G)$ is a $k$-dominating if for every $u\in V(G)$, there exists a path of length at most $k$ linking $u$ and $X$.
Throughout the article, we use the notion of treewidth. However, for the understanding of the current article, the exact definition of treewidth, which can for example be found in [4], is not required. The properties described in Proposition 2.2 and Theorem 2.3 as well as the well-known fact that the treewidth of a graph is the maximum of the treewidth of one of its components will be sufficient.
Recall that an instance of Multicut consists of a graph $G$ and a graph $H$ with $V(H)\subseteq V(G)$ and a multicut for $(G,H)$ is a set $S\subseteq E(G)$ such that $t_{1}$ and $t_{2}$ are in distinct components of $G\setminus S$ for all $t_{1},t_{2}\in V(H)$ with $t_{1}t_{2}\in E(H)$. For the edge-weighted version of the problem, we assume that the input graph $G$ is equipped with a weight function $w:E(G)\to\mathbb{Z}^{+}$.
Appendix B Further possible parameterizations
We here mention two further notions of almost planar graphs that we are not studying in this paper.
First,t would be very natural to consider graphs that can be made planar by the deletion of $\pi$ vertices. However, the problem can be very hard even for small values of this parameter. Specifically, the Multiway Cut problem with three terminals is $\mathsf{NP}$-hard even for graphs that can be made planar by the deletion of three vertices, as the Group 3-Terminal Cut problem on planar graphs can be reduced to such an instance [5]. Thus we cannot expect results similar to Theorems 1.21.8 for this parameter.
Further, a graph is 1-planar if it can be drawn in the plance such that every edge participates in at most one crossing. Every graph can be made 1-planar by subdividing each edge a sufficient number of times. Subdividing an edge (and giving the same weight to both new edges) does not change the Multicut problem, hence the problem is as hard on 1-planar graphs as on general graphs. Therefore, 1-planar graphs are irrelevant for the study of Multicut and its special cases.
Appendix C Proof of Theorem 1.5
The objective of this section is to prove Theorem 1.5, which we restate here for convenience. See 1.5
Roughly speaking, the remainder of Section C can be summarized as follows. First, in Section C.1, we collect some simple preliminary results for our algorithm. Next, in Section C.2, we give some structural results on minimum multicuts we need for our algorithm. In particular, we prove Lemma 3.2. Next, in Section C.3, we prove that arbitrary states can efficiently be made relevant, that is, we prove Lemma 3.5. After, in Section C.4, we show how a relevant state can be replaced by a small number of states whose value for $\kappa$ is larger, meaning we prove Lemma 3.6. Next, in Section C.5, we show how to deal with complete states. More precisely, we prove Lemmas 3.3 and 3.4 that show how to obtain Lemma 3.1 once a complete valid state is available In Section C.6, we prove Lemma 3.1 from the results obtained in the previous sections. Finally, in Section C.7, we conclude Theorem 1.5 from Lemma 3.1.
C.1 Preparatory results
In this section, we collect some preliminary results we need for our algorithm. First, a solution of the multicut problem can be checked in polynomial time (e.g., with the use of breadth first search).
Proposition C.1.
Let $(G,H)$ be an instance of Multicut and $S\subseteq E(G)$. Then we can check in polynomial time whether $S$ is a multicut for $(G,H)$.
We next need the following simple property which holds for optimization problems aiming for the minimization of the number of elements we delete in general. Its proof can be found in [7].
Proposition C.2.
Let $(G,H)$ be an instance of Multicut with $S$ being a minimum multicut for $(G,H)$ and $F\subseteq S$. Then for every minimum multicut $S^{\prime}$ for $(G\setminus F,H)$, we have that $S^{\prime}\cup F$ is a minimum multicut for $(G,H)$.
The next two results we need deal with the number of edges that are needed to make a given graph planar. We first need the following result, showing that a small edge set whose deletion makes a graph planar can be found sufficiently fast for our purposes. Analogous results are known also for the vertex-deletion version [15, 1, 11], which can be made to work also for the edge-deletion version by appropriate reductions.
Proposition C.3 ([12]).
Let $G$ be a graph and $\pi$ a nonnegative integer. Then there exists an algorithm that either outputs a set $E_{\pi}\subseteq E(G)$ such that $|E_{\pi}|\leq\pi$ and $G\setminus E_{\pi}$ is planar or correctly reports that no such set exists and runs in $f(\pi)n$.
The next lemma is important as it shows that progress is made on an instance when an edge linking two connected components of the planar part of the graph is detected. Its proof can be found in [7].
Lemma C.4.
Let $G$ be a graph and $E^{\prime}\subseteq E(G)$ be such that $G\setminus E^{\prime}$ is planar. If there is an edge $e$ in $E^{\prime}$ whose endvertices are contained in distinct components of $G\setminus E^{\prime}$, then $G\setminus(E^{\prime}\setminus\{e\})$ is planar.
Further, we give some basic results on relevant sets. For the first of these results, the uniqueness of $Y_{3}$ is well-known and can be established by a simple submodularity argument. Moreover, computing $Y_{3}$ is possible through a simple modification of the algorithm of Edmonds and Karp and the last property follows directly from the minimality of $Y_{3}$. Propositions C.6 and C.7 follow directly from the definition of relevant sets.
Proposition C.5.
Let $G$ be a connected graph and $Y_{1},Y_{2}\subseteq V(G)$ be disjoint sets. Then there exists a unique relevant set $Y_{3}$ for $(Y_{1},Y_{2})$. Moreover $Y_{3}$ can be computed in polynomial time and every component of $G[Y_{3}]$ contains a component of $G[Y_{1}]$.
Proposition C.6.
Let $G$ be a graph and $Y_{1},Y_{2}\subseteq V(G)$ be disjoint, $Y_{3}$ the relevant set for $(Y_{1},Y_{2})$ and $Y_{4}\subseteq V(G)$ with $Y_{1}\subseteq Y_{4}\subseteq Y_{3}$. Then $Y_{3}$ is the relevant set for $(Y_{4},Y_{2})$.
Proposition C.7.
Let $G$ be a graph and $Y_{1},Y_{2}\subseteq V(G)$ be disjoint, $Y_{3}$ the relevant set for $(Y_{1},Y_{2})$ and $Y_{4}\subseteq V(G)$ with $Y_{2}\subseteq Y_{4}\subseteq V(G)\setminus Y_{3}$. Then $Y_{3}$ is the relevant set for $(Y_{1},Y_{4})$.
Throughout our algorithm, it is important to assure that the value of $\kappa$ increases continuously. The following easy result establishes a criterion to test whether this is actually the case.
Proposition C.8.
Let $(G,H)$ be a $\textup{\sf planar}+\pi e$ instance of Multicut, and $\mathcal{P}^{1}$ and $\mathcal{P}^{2}$ two states for $(G,H)$ such that $\mathcal{P}^{2}$ extends $\mathcal{P}^{1}$. Then $\kappa(\mathcal{P}^{2})\geq\kappa(\mathcal{P}^{1})$. Moreover, if there exists some $Y_{0}\in\mathcal{P}^{1}$ and $\overline{Y_{0}}\in\mathcal{P}^{2}$ with $Y_{0}\subseteq\overline{Y_{0}}$ and $\kappa(\mathcal{P}^{1},Y_{0})<\kappa(\mathcal{P}^{2},\overline{Y_{0}})$, then $\kappa(\mathcal{P}^{2})>\kappa(\mathcal{P}^{1})$.
Proof.
As $\mathcal{P}^{2}$ extends $\mathcal{P}^{1}$, for every $Y\in\mathcal{P}^{1}$, there exists a unique $\overline{Y}\in\mathcal{P}^{2}$ such that $Y\subseteq\overline{Y}$ with $\overline{Y_{1}}\neq\overline{Y_{2}}$ for all distinct $Y_{1},Y_{2}\in\mathcal{P}^{1}$.
We first show that $\alpha_{G_{0},\mathcal{P}^{1}}(Y)\leq\alpha_{G_{0},\mathcal{P}^{2}}(\overline{Y})$ for all $Y\in\mathcal{P}$. Consider some $Y\in\mathcal{P}$. As $Y\subseteq\overline{Y}$ and $\bigcup\mathcal{P}^{1}\setminus Y\subseteq\bigcup\mathcal{P}^{2}\setminus\overline{Y}$, we obtain that for every set $S\subseteq E(G_{0})$ whose deletion separates all vertices in $Y$ from all vertices in $\bigcup\mathcal{P}^{1}\setminus Y$, we have that its deletion also separates all vertices in $\overline{Y}$ from all vertices in $\bigcup\mathcal{P}^{2}\setminus\overline{Y}$. This yields $\lambda_{G_{0}}(Y,\bigcup\mathcal{P}^{1}\setminus Y)\leq\lambda_{G_{0}}(\overline{Y},\bigcup\mathcal{P}^{2}\setminus\overline{Y})$. As $\mathcal{P}^{1}$ and $\mathcal{P}^{2}$ are states, we obtain that $\alpha_{G_{0},\mathcal{P}^{1}}(Y)-\alpha_{G_{0},\mathcal{P}^{2}}(\overline{Y})=\lambda_{G_{0}}(Y,\bigcup\mathcal{P}^{1}\setminus Y)-\lambda_{G_{0}}(\overline{Y},\bigcup\mathcal{P}^{2}\setminus\overline{Y})\leq 0$.
We now show that $\kappa(\mathcal{P}^{1},Y)\leq\kappa(\mathcal{P}^{2},\overline{Y})$ for all $Y\in\mathcal{P}^{1}$. Consider some $Y\in\mathcal{P}$. We distinguish some cases on the shape of $Y$. First suppose that $Y\in\mathcal{P}^{1}_{\textup{fat}}$. Then we have $\alpha_{G_{0},\mathcal{P}^{2}}(\overline{Y})\geq\alpha_{G_{0},\mathcal{P}^{1}}(Y)\geq\pi(t+1)+1$, so $\overline{Y}\in\mathcal{P}^{2}_{\textup{fat}}$. It follows that $\kappa(\mathcal{P}^{2}.\overline{Y})=2(\pi(t+1)+1)=\kappa(\mathcal{P}^{1},Y)$. Next suppose that $Y\in\mathcal{P}^{1}_{\textup{f-n}}$, so there exists some $Y_{0}\in\mathcal{P}^{1}_{\textup{fat}},u\in Y$, and $v\in Y_{0}$ such that $uv\in E(G_{0})$. By the same argument as before, we obtain that $\overline{Y_{0}}\in\mathcal{P}^{2}_{\textup{fat}}$. It follows that $\overline{Y}\in\mathcal{P}^{2}_{\textup{fat}}\cup\mathcal{P}^{2}_{\textup{f-n}}$. If $\overline{Y}\in\mathcal{P}^{2}_{\textup{fat}}$, we obtain that $\kappa(\mathcal{P}^{2},\overline{Y})=2(\pi(t+1)+1)>\pi(t+1)+1+\alpha_{G_{0},\mathcal{P}^{1}}(Y)=\kappa(\mathcal{P}^{1},Y)$. If $\overline{Y}\in\mathcal{P}^{2}_{\textup{f-n}}$, we have $\kappa(\mathcal{P}^{2},\overline{Y})-\kappa(\mathcal{P}^{1},Y)=\alpha_{G_{0},\mathcal{P}^{2}}(\overline{Y})-\alpha_{G_{0},\mathcal{P}^{1}}(Y)\geq 0$. Finally suppose that $Y\in\mathcal{P}^{1}_{\textup{thin}}$. If $\overline{Y}\in\mathcal{P}^{2}_{\textup{fat}}\cup\mathcal{P}^{2}_{\textup{f-n}}$, we obtain $\kappa(\mathcal{P}^{2},\overline{Y})\geq\pi(t+1)+1>\alpha_{G_{0},\mathcal{P}^{1}}(Y)=\kappa(\mathcal{P}^{1},Y)$. If $\overline{Y}\in\mathcal{P}^{2}_{\textup{thin}}$, we have $\kappa(\mathcal{P}^{2},\overline{Y})-\kappa(\mathcal{P}^{1},Y)=\alpha_{G_{0},\mathcal{P}^{2}}(\overline{Y})-\alpha_{G_{0},\mathcal{P}^{1}}(Y)\geq 0$.
We obtain $\kappa(\mathcal{P}^{1})=\sum_{Y\in\mathcal{P}^{1}}\kappa(\mathcal{P}^{1},Y)\leq\sum_{Y\in\mathcal{P}^{1}}\kappa(\mathcal{P}^{2},\overline{Y})\leq\kappa(\mathcal{P}^{2})$. Moreover, if equality holds, then $\kappa(\mathcal{P}^{1},Y)=\kappa(\mathcal{P}^{2},\overline{Y})$ holds for every $Y\in\mathcal{P}^{1}$. ∎
C.2 Structural properties of minimum multicuts
This section is dedicated to proving some structural properties of minimum multicuts, which are crucially needed in our algorithm. In particular, we prove Lemma 3.2. The first result is simple, but useful.
Proposition C.9.
Let $(G,H)$ be a connected $\textup{\sf planar}+\pi e$ instance, $S$ a minimum multicut for $(G,H)$ and $Q$ a component of $G_{0}\setminus S_{0}$. Then either $V(Q)\cap T\neq\emptyset$ or there exists some $e\in E_{\pi}$ that has exactly one endvertex in $V(Q)$.
Proof.
Suppose otherwise. Observe that $Q$ is a component of $G\setminus S$. As $G$ is connected and $\delta_{G}(V(Q))\cap E_{\pi}=\emptyset$ by assumption, there exists some $e\in\delta_{G_{0}}(V(Q))$. Let $Q_{0}$ be the unique component of $G\setminus S$ containing the unique endvertex of $e$ that is not contained in $V(Q)$. Let $S^{\prime}=S\setminus\{e\}$. Let $t_{1},t_{2}\in T$ with $t_{1}t_{2}\in E(H)$ and for $i\in[2]$, let $Q_{i}$ be the unique component of $G\setminus S$ containing $t_{i}$. Observe that $Q\notin\{Q_{1},Q_{2}\}$ by the assumption that $V(Q)\cap T=\emptyset$. Further, as $S$ is a multicut for $(G,H)$, we have that $Q_{1}\neq Q_{2}$. Hence, by symmetry, we may suppose that $Q_{0}\neq Q_{1}$. We obtain that $\delta_{G}(V(Q_{1}))\subseteq S^{\prime}$. It follows that $t_{1}$ and $t_{2}$ are in distinct components of $G\setminus S^{\prime}$. This contradicts the minimality of $S$. ∎
The objective of the remainder of this section is to prove Lemma 3.2. We first need the following preparatory result.
Lemma C.10.
Let $(G,H)$ be a $\textup{\sf planar}+\pi e$ instance of Multicut, let $S$ be a minimum multicut for $(G,H)$, let $i\in[2]$, and let $Q_{1},Q_{2}$ be two components of $G\setminus S_{0}$ such that $(V(Q_{1})\cup V(Q_{2}))\cap T\subseteq B_{i}\cup I$. Then $\delta_{G_{0}}(V(Q_{1}),V(Q_{2}))\leq\pi$.
Proof.
Suppose otherwise and let $S^{\prime}=(S\setminus\delta_{G_{0}}(V(Q_{1}),V(Q_{2})))\cup E_{\pi}$. As $\delta_{G_{0}}(V(Q_{1}),V(Q_{2}))\subseteq S$ and $|E_{\pi}|=\pi$ by assumption, we have $|S^{\prime}|<|S|$.
We next show that $S^{\prime}$ is a multicut for $(G,H)$. To this end, let $t_{1},t_{2}\in T$ with $t_{1}t_{2}\in E(H)$. As $(B_{1},B_{2},I,X)$ is an extended biclique decomposition for $(G,H)$ and by symmetry, we may suppose that $t_{1}\in B_{3-i}\cup X$. Let $Q$ be the component of $G_{0}\setminus S_{0}$ containing $t_{1}$ and let $Q^{\prime}$ be the component of $G\setminus S^{\prime}$ containing $t_{1}$. Observe that by construction, we have $Q\notin\{Q_{1},Q_{2}\}$ and hence $\delta_{G_{0}}(V(Q))\cup E_{\pi}\subseteq S^{\prime}$, so $V(Q^{\prime})\subseteq V(Q)$. As $S$ is a multicut for $(G,H)$, we have that $t_{2}\notin V(Q)$ and hence $t_{2}\notin V(Q^{\prime})$. It follows that $S^{\prime}$ is a multicut for $(G,H)$. As $|S^{\prime}|<|S|$, this contradicts $S$ being a minimum multicut for $(G,H)$. ∎
We are now ready to prove Lemma 3.2, which we restate here for convenience. See 3.2
Proof.
Suppose otherwise, so $\textup{dist}_{G_{0},S_{0}}(Y,X)\geq 2$. By assumption, we in particular have $Y\cap X=\emptyset$. Hence, as $S$ is a multicut for $(G,H)$ and by symmetry, we may suppose that $Y\cap T\subseteq B_{1}\cup I$. Next, let $\mathcal{Q}_{1}$ be the set of components of $G_{0}\setminus S_{0}$ that are linked to $Y$ by at least one edge and that contain at least one element of $B_{1}$ and let $F_{1}\subseteq E(G_{0})$ be the set of edges in $E(G_{0})$ connecting $Y$ and $\bigcup_{Q\in\mathcal{Q}_{1}}V(Q)$. As $\textup{dist}_{G_{0},S_{0}}(Y,X)\geq 2$, we have that $V(Q)\cap X=\emptyset$ for all $Q\in\mathcal{Q}_{1}$. It hence follows by Lemma C.10 that $d_{G_{0}}(Y,V(Q))\leq\pi$ for all $Q\in\mathcal{Q}_{1}$. As we clearly have $|\mathcal{Q}_{1}|\leq t$, we obtain $|F_{1}|\leq\pi t$. Now let $R\subseteq V(G_{0})$ be a set such that $\delta_{G_{0}}(R)$ is a $(Y\cap(W\cup T),(W\cup T)\setminus Y)$-cut of size $\lambda(Y\cap(W\cup T),(W\cup T)\setminus Y)$ in $G_{0}$. We now define $S^{\prime}=(S\setminus\delta_{G_{0}}(Y))\cup(\delta_{G_{0}}(R)\cup F_{1}\cup E_{\pi})$. As $\delta_{G_{0}}(Y)\subseteq S$, by the definitions of $R$ and $F_{1}$ and as $Y$ is fat in $\mathcal{P}$, we have
$|S^{\prime}|$ $\leq|S|-d_{G_{0}}(Y)+d_{G_{0}}(R)+|F_{1}|+|E_{\pi}|$
$\leq|S|-(d_{G_{0}}(Y)-\lambda(Y\cap(W\cup T),(W\cup T)\setminus Y))+\pi t+\pi$
$=|S|-\alpha_{G_{0},\mathcal{P}}(Y)+\pi(t+1)$
$<|S|.$
We next show that $S^{\prime}$ is a multicut for $(G,H)$. Let $t_{1},t_{2}\in T$ with $t_{1}t_{2}\in E(H)$. First suppose that $\{t_{1},t_{2}\}\cap R\neq\emptyset$, say $t_{1}\in R$. As $S$ is a multicut for $(G,H)$, it follows that $t_{2}\notin Y$, so in particular $t_{2}\notin R$. As $\delta_{G_{0}}(R)\cup E_{\pi}\subseteq S^{\prime}$, we obtain that $t_{1}$ and $t_{2}$ are in distinct components of $G\setminus S^{\prime}$.
Now suppose that $\{t_{1},t_{2}\}\cap R=\emptyset$. As $(B_{1},B_{2},I,X)$ is an extended biclique decoposition of $(G,H)$ and by symmetry, we may suppose that $t_{1}\in B_{1}\cup X$. Let $Q$ be the component of $G_{0}\setminus S_{0}$ containing $t_{1}$ and let $Q^{\prime}$ be the component of $G\setminus S^{\prime}$ containing $t_{1}$. As $\textup{dist}_{G_{0},S_{0}}(Y,X)\geq 2$ and by construction, we obtain that either $Q\in\mathcal{Q}_{1}$ or $E(G_{0})$ does not contain an edge linking $V(Q)$ and $Y$. In either case, we obtain that $\delta_{G_{0}}(V(Q))\cup E_{\pi}\subseteq S^{\prime}$. It follows that $V(Q^{\prime})\subseteq V(Q)$. As $S$ is a multicut for $(G,H)$, we obtain that $t_{2}\notin V(Q^{\prime})$. Hence $S^{\prime}$ is a multicut for $(G,H)$. As $|S^{\prime}|<|S|$, this contradicts $S$ being a minimum multicut for $(G,H)$. ∎
C.3 Making states relevant
This section is dedicated to showing that, given an arbitrary state $\mathcal{P}$, we can efficiently compute another state $\mathcal{P}^{\prime}$, which is relevant and maintains all important properties of $\mathcal{P}$. In other words, we prove Lemma 3.5. Most technical difficulties for this proof are contained in the following result, which shows that a single class of a state can be replaced by a larger one which is relevant in the arising state without losing any important property.
Lemma C.11.
Let $(G,H)$ be a connected $\textup{\sf planar}+\pi e$ instance of Multicut, and let $\mathcal{P}$ be a state. Further, let $Y_{0}\in\mathcal{P}$, let $Y_{0}^{\prime}$ be the relevant set for $(Y_{0},\bigcup\mathcal{P}\setminus Y_{0})$ in $G_{0}$ and let $\mathcal{P}^{\prime}$ be obtained from $\mathcal{P}$ by replacing $Y_{0}$ by $Y_{0}^{\prime}$. Then $\mathcal{P}^{\prime}$ is a state with $\tau(\mathcal{P}^{\prime})\leq\tau(\mathcal{P})$ and $\kappa(\mathcal{P}^{\prime})\geq\kappa(\mathcal{P})$ and if $\mathcal{P}$ is maximum valid, so is $\mathcal{P}^{\prime}$.
Proof.
It follows directly by construction that $\mathcal{P}^{\prime}$ is a state. We now show that $\tau(\mathcal{P}^{\prime})\leq\tau(\mathcal{P})$ and $\kappa(\mathcal{P}^{\prime})\geq\kappa(\mathcal{P})$. By Proposition C.5, we have that every component of $Y_{0}^{\prime}$ contains a component of $Y_{0}$. By the definition of $\mathcal{P}^{\prime}$, this yields $\tau(\mathcal{P}^{\prime})-\tau(\mathcal{P})=\tau(\{Y_{0}^{\prime}\})-\tau(\{Y_{0}\})\leq 0$.
Next observe that $\kappa(\mathcal{P}^{\prime})\geq\kappa(\mathcal{P})$ holds by Proposition C.8.
For the last part, we assume that $\mathcal{P}$ is maximum valid. We need to show that $\mathcal{P}^{\prime}$ is maximum valid. As $\mathcal{P}$ is maximum valid, there exists a minimum multicut $S$ for $(G,H)$ that respects $\mathcal{P}$. It follows in particular that for every $Y\in\mathcal{P}$, there exists some $\overline{Y}\in\mathcal{K}(G_{0}\setminus S_{0})$ with $Y\subseteq\overline{Y}$. As $\mathcal{K}(G_{0}\setminus S_{0})$ extends $\mathcal{P}$ and the maximality of $\mathcal{P}$, we obtain that $\mathcal{K}(G_{0}\setminus S_{0})=\{\overline{Y}:Y\in\mathcal{P}\}$. We define $\tilde{Y_{0}}=\overline{Y_{0}}\cup Y_{0}^{\prime}$ and $\tilde{Y}=\overline{Y}\setminus Y_{0}^{\prime}$ for all $Y\in\mathcal{P}\setminus\{Y_{0}\}$. Observe that, as $W\cup T\subseteq\bigcup\mathcal{P}$, as $\mathcal{K}(G_{0}\setminus S_{0})$ extends $\mathcal{P}$, and by the definition of $Y_{0}^{\prime}$, we have $(W\cup T)\cap Y=(W\cup T)\cap\overline{Y}=(W\cup T)\cap\tilde{Y}$ for all $Y\in\mathcal{P}$. In particular, $\mathcal{P}^{\prime}$ is a state. Next, we set $S^{\prime}=(S\setminus\delta_{G_{0}}(\overline{Y_{0}}))\cup\delta_{G_{0}}(\tilde{Y_{0}})$. We will show that $S^{\prime}$ is a minimum multicut for $(G,H)$ and $\mathcal{K}(G_{0}\setminus S^{\prime}_{0})=\{\tilde{Y}:Y\in\mathcal{P}\}$. This directly implies that $S^{\prime}$ respects $\mathcal{P}^{\prime}$.
The following result is a first step in the direction of characterizing the components of $G_{0}\setminus S_{0}^{\prime}$.
Claim C.12.
For every $Y\in\mathcal{P}$, we have that $\tilde{Y}$ is the union of a collection of classes of $\mathcal{K}(G_{0}\setminus S^{\prime}_{0})$.
Proof.
[…]
We are now ready to show that $S^{\prime}$ is a multicut indeed.
Claim C.13.
$S^{\prime}$ is a multicut for $(G,H)$.
Proof.
[…]
We next show that $S^{\prime}$ is of minimum size.
Claim C.14.
$|S^{\prime}|\leq|S|$.
Proof.
As $Y_{0}^{\prime}$ is the relevant set for $(Y_{0},\bigcup\mathcal{P}\setminus Y_{0})$ in $G_{0}$, we obtain that $d_{G_{0}}(\overline{Y_{0}}\cap Y_{0}^{\prime})\geq d_{G_{0}}(Y_{0}^{\prime})$. By the submodularity of the degree function, we obtain that $d_{G_{0}}(\tilde{Y_{0}})\leq d_{G_{0}}(\overline{Y_{0}})+d_{G_{0}}(Y_{0}^{\prime})-d_{G_{0}}(\overline{Y_{0}}\cap Y_{0}^{\prime})\leq d_{G_{0}}(\overline{Y_{0}})$.
By the minimality of $S$, together with $\delta_{G_{0}}(\overline{Y_{0}})\subseteq S$, this yields
$|S^{\prime}|$ $=|(S\setminus\delta_{G_{0}}(\overline{Y_{0}}))\cup\delta_{G_{0}}(\tilde{Y_{0}})|$
$\leq|S\setminus\delta_{G_{0}}(\overline{Y_{0}})|+d_{G_{0}}(\tilde{Y_{0}})$
$=|S|-d_{G_{0}}(\overline{Y_{0}})+d_{G_{0}}(\tilde{Y_{0}})$
$\leq|S|.$
[…]
It follows from Claims C.13 and C.14 and the choice of $S$ that $S^{\prime}$ is a minimum multicut for $(G,H)$. With this result at hand, we are able to use the maximality of $\mathcal{P}$ to fully characterize the components of $G_{0}\setminus S^{\prime}_{0}$. The next claim will directly imply that $\mathcal{K}(G_{0}\setminus S^{\prime}_{0})$ extends $\mathcal{P}^{\prime}$ and hence finish the proof.
Claim C.15.
$\mathcal{K}(G_{0}\setminus S^{\prime}_{0})=\{\tilde{Y}:Y\in\mathcal{P}\}$.
Proof.
[…]
We are now ready to prove Lemma 3.5, which we restate here for convenience. See 3.5
Proof.
Let $Y_{1},\ldots,Y_{q}$ be an arbitrary enumeration of $\mathcal{P}$. We now define a sequence $\mathcal{P}^{0},\ldots,\mathcal{P}^{q}$ of states extending $\mathcal{P}$ such that $Y_{i+1},\dots,Y_{q}\in\mathcal{P}^{i}$ for $i\in[q]$. First, we set $\mathcal{P}^{0}=\mathcal{P}$. Next, for $i\in[q]$, let $Y_{i}^{\prime}$ be the relevant set for $(Y_{i},\bigcup\mathcal{P}^{i-1}\setminus Y_{i})$. We obtain $\mathcal{P}^{i}$ from $\mathcal{P}^{i-1}$ by replacing $Y_{i}$ by $Y^{\prime}_{i}$. We set $\mathcal{P}^{\prime}=\mathcal{P}^{q}$. It follows by construction and Proposition C.7 that $Y^{\prime}$ is the relevant set for $(Y^{\prime},\bigcup\mathcal{P}^{\prime}\setminus Y^{\prime})$ for all $Y\in\mathcal{P}$. Hence $\mathcal{P}^{\prime}$ is a relevant state. Moreover, repeatedly applying Lemma C.11, we obtain that $\tau(\mathcal{P}^{\prime})\leq\tau(\mathcal{P})$, $\kappa(\mathcal{P}^{\prime})\geq\kappa(\mathcal{P})$ and that, if $\mathcal{P}$ is maximum valid, then so is $\mathcal{P}^{\prime}$. Further, we have $|\mathcal{P^{\prime}}|=|\mathcal{P}|$ by construction. Finally, it follows from Proposition C.5 and that fact $q\leq\pi+t$ that $\mathcal{P}^{\prime}$ can be computed in polynomial time. ∎
C.4 Dealing with incomplete states
In this section, we show how to replace an incomplete state by a bounded number of states for which $\kappa$ is larger. More precisely, we prove Lemma 3.6, which we restate here for convenience. See 3.6
Proof.
As $\mathcal{P}$ is incomplete, there exists a vertex $u_{0}\in V(G)\setminus\bigcup\mathcal{P}$ that has a neighbor in some $Y_{0}\in\mathcal{P}_{\textup{thin}}$ in $G_{0}$. Now for every $Y\in\mathcal{P}$, let $Y^{\prime}=Y\cup\{u_{0}\}$ and let $\mathcal{P}^{Y}$ be the subpartition of $V(G)$ in which $Y$ is replaced by $Y^{\prime}$. It is easy to see that $\mathcal{P}^{Y}$ is a state for every $Y\in\mathcal{P}$. We now show that $(\mathcal{P}^{Y})_{Y\in\mathcal{P}}$ has the desired properties. First observe that $(\mathcal{P}^{Y})_{Y\in\mathcal{P}}$ can be computed in polynomial time by construction. Next, we have $|(\mathcal{P}^{Y})_{Y\in\mathcal{P}}|=|\mathcal{P}|\leq\pi+t$ as $\mathcal{P}$ is a state.
Let $Y\in\mathcal{P}$. Then $\mathcal{P}^{Y}$ is obtained from $\mathcal{P}$ by replacing $Y$ with $Y^{\prime}$. As $Y^{\prime}=Y\cup\{u_{0}\}$, we obtain that $G_{0}[Y^{\prime}]$ has at most one more component than $G_{0}[Y]$. This yields $\tau(\mathcal{P}^{Y})\leq\tau(\mathcal{P})+1$.
We next show that $\kappa$ is increased for all newly created states, which is slightly more technical.
Claim C.16.
$\kappa(\mathcal{P}^{Y})>\kappa(\mathcal{P})$ for all $Y\in\mathcal{P}$.
Proof.
Let $Y\in\mathcal{P}$. First observe that $\mathcal{P}^{Y}$ clearly extends $\mathcal{P}$.
First consider the case that $Y\in\mathcal{P}_{\textup{f-n}}\cup\mathcal{P}_{\textup{thin}}$. We first show that $\kappa(\mathcal{P}^{Y},Y^{\prime})>\kappa(\mathcal{P},Y)$. If $Y^{\prime}\in\mathcal{P}^{Y}_{\textup{fat}}$, we obtain that $\kappa(\mathcal{P}^{Y},Y^{\prime})-\kappa(\mathcal{P},Y)\geq 2(\pi(t+1)+1)-(\pi(t+1)+1+\alpha_{G_{0},\mathcal{P}}(Y))=\pi(t+1)+1-\alpha_{G_{0},\mathcal{P}}(Y)>0$. We may hence suppose that $Y^{\prime}\in\mathcal{P}^{Y}_{\textup{f-n}}\cup\mathcal{P}^{Y}_{\textup{thin}}$. It follows that $\kappa(\mathcal{P}^{Y},Y^{\prime})-\kappa(\mathcal{P},Y)\geq\alpha_{G_{0},\mathcal{P}^{Y}}(Y^{\prime})-\alpha_{G_{0},\mathcal{P}}(Y)$. As $\mathcal{P}$ is relevant, we have that $Y$ is the relevant set for $(Y,\bigcup\mathcal{P}\setminus Y)$, which yields $\lambda(Y^{\prime},\bigcup\mathcal{P}\setminus Y)>\lambda(Y,\bigcup\mathcal{P}\setminus Y)$. By construction, we have $\alpha_{G_{0},\mathcal{P}^{Y}}(Y^{\prime})-\alpha_{G_{0},\mathcal{P}}(Y)=\lambda(Y^{\prime},\bigcup\mathcal{P}^{Y}\setminus Y^{\prime})-\lambda(Y,\bigcup\mathcal{P}\setminus Y)=\lambda(Y^{\prime},\bigcup\mathcal{P}\setminus Y)-\lambda(Y,\bigcup\mathcal{P}\setminus Y)>0$. In either case, we obtain $\kappa(\mathcal{P}^{Y},Y^{\prime})>\kappa(\mathcal{P},Y)$. It follows that $\kappa(\mathcal{P}^{Y})>\kappa(\mathcal{P})$ by Proposition C.8.
[…]
For the rest of the proof, suppose that $\mathcal{P}$ is maximum valid, so there exists a minimum multicut $S$ for $(G,H)$ that respects $\mathcal{P}$. It follows that for every $Y\in\mathcal{P}$, there exists a unique $\overline{Y}\in\mathcal{K}(G_{0}\setminus S_{0})$ such that $Y\subseteq\overline{Y}$ with $\overline{Y_{1}}\neq\overline{Y_{2}}$ for all distinct $Y_{1},Y_{2}\in\mathcal{P}$. By the maximality of $\mathcal{P}$, we obtain that $\mathcal{K}(G_{0}\setminus S_{0})=(\overline{Y}:Y\in\mathcal{P})$. In particular, there exists a unique $Y_{1}\in\mathcal{P}$ such that $u_{0}\in\overline{Y_{1}}$. It follows by definition that $S$ respects $\mathcal{P}^{Y_{1}}$, so $\mathcal{P}^{Y_{1}}$ is valid. As $|\mathcal{P}^{Y_{1}}|=|\mathcal{P}|$ and $\mathcal{P}$ is maximum valid, so is $\mathcal{P}^{Y_{1}}$. ∎
C.5 Dealing with complete states
In this section, we show how to continue with our algorithm once a complete maximum valid state $\mathcal{P}$ is attained. More precisely, we prove Lemma 3.3 and Lemma 3.4. The case that the state in consideration contains a thin class is dealt with in Lemma 3.3. The other case is when $\mathcal{P}_{\textup{thin}}$ is empty. In this case, we show that a minimum multicut can be directly computed. We need to combine some topological arguments with the machinery developped in [7]. The result is proved in Lemma 3.4.
We start by dealing with the first case. More precisely, we prove Lemma 3.3, which we restate here for convenience. See 3.3
Proof.
We first choose some thin $Y_{0}\in\mathcal{P}$ and define $S^{\prime}=\delta_{G_{0}}(Y_{0})$. Clearly, $S^{\prime}$ can be computed in polynomial time.
Next, as $\mathcal{P}$ is valid, there exists a minimum multicut $S$ for $(G,H)$ that respects $\mathcal{P}$. For $Y\in\mathcal{P}$, as $\mathcal{K}(G_{0}\setminus S_{0})$ extends $\mathcal{P}$, there exists a unique $\overline{Y}\in\mathcal{K}(G_{0}\setminus S_{0})$ with $Y\subseteq\overline{Y}$. Now consider an edge $v_{0}v_{1}\in S^{\prime}$ with $v_{0}\in Y_{0}$. As $\mathcal{P}$ is complete, there exists some $Y_{1}\in\mathcal{P}$ such that $v_{1}\in Y_{1}$. We obtain that $v_{0}\in\overline{Y_{0}}$ and $v_{1}\in\overline{Y_{1}}$. It follows that $S^{\prime}\subseteq S$ and that $\overline{Y_{0}}=Y_{0}$. For the remainder of the proof, we need a small case distinction. By Proposition C.9 and the fact that $Y_{0}\in\mathcal{K}(G_{0}\setminus S_{0})$, one of the following two cases occurs. Further, we can clearly observe in polynomial time which of the two cases occurs.
Case C.17.
There exists an edge $e\in E_{\pi}$ with exactly one endvertex in $Y_{0}$.
Proof.
[…]
Case C.18.
$Y_{0}\cap W=V(E^{\prime})$ for some $E^{\prime}\subseteq E_{\pi}$ and $Y\cap T\neq\emptyset$.
Proof.
We now define $G^{\prime}=G-Y_{0}$ and $H^{\prime}=H-Y_{0}$. Clearly, $(G^{\prime},H^{\prime})$ can be computed in polynomial time and is an extended subinstance of $(G,H)$. Further, $|V(H^{\prime})|<|V(H)|$ holds by assumption.
Now let $S^{\prime\prime}$ be a minimum multicut for $(G^{\prime},H^{\prime})$. We first show that $S^{\prime}\cup S^{\prime\prime}$ is a multicut for $(G,H)$. Let $t_{1},t_{2}\in V(H)$ with $t_{1}t_{2}\in E(H)$. If $\{t_{1},t_{2}\}\subseteq Y_{0}$, we obtain an immediate contradiction as $Y_{0}$ is a component of $G\setminus S$ and $S$ is a multicut for $(G,H)$. We may hence suppose that $\{t_{1},t_{2}\}\setminus Y_{0}\neq\emptyset$. Suppose for the sake of a contradiction that there exists a $t_{1}t_{2}$-path $P$ in $G\setminus(S^{\prime}\cup S^{\prime\prime})$. If $V(P)\cap Y_{0}=\emptyset$, we obtain that $P$ also exists in $G^{\prime}\setminus S^{\prime\prime}$, a contradiction to $S^{\prime\prime}$ being a multicut for $(G^{\prime},H^{\prime})$. Otherwise, we obtain that $E(P)$ contains an edge $e$ with exactly one endvertex in $Y_{0}$. If $e\in E(G_{0})$, we obtain a contradiction as $\delta_{G_{0}}(Y_{0})\subseteq S^{\prime}\cup S^{\prime\prime}$. If $e\in E_{\pi}$, we obtain a contradiction to the assumption that $Y_{0}\cap W=V(E^{\prime})$ for some $E^{\prime}\subseteq E_{\pi}$.
[…]
We now consider the case that $\mathcal{P}$ does not contain a thin set. Our objective is to prove Lemma 3.4. We want to make use of the fact that the problem can be reduced to computing a multicut of a planar instance of Multicut such that every minimum multicut dual for this instance is of small treewidth. The key idea is that for any minimum multicut $S$ for $(G,H)$, we have that $\textup{dist}_{G_{0},S_{0}}(v,X)$ is small for all $v\in V(G)$. The first result shows that this property indeed guarantees that all minimum multicut duals are of small treewidth.
Lemma C.19.
Let $(G,H)$ be a planar instance of Multicut, let $(B_{1},B_{2},I,X)$ an extended biclique decomposition of $H$ with $|X|=\mu$, let $C$ be an inclusionwise minimal multicut dual for $(G,H)$, let $S=e_{G}(C)$ and suppose that $\textup{dist}_{G,S}(v,X)\leq 2$ for every $v\in V(G)$. Then $\textup{tw}(C)=O(\sqrt{\mu})$.
Proof.
We now construct a graph $C^{\prime}$ from $C$ by adding a new vertex $z_{f}$ for every face $f$ of $C$ and linking it to all vertices in $V(C)$ incident to $f$. Observe that $C^{\prime}$ is planar. Next, let $Z_{X}\subseteq V(C^{\prime})$ be the set containing the vertex $z_{f}$ for every face $f$ of $C$ containing some $x\in X$.
Claim C.20.
$Z_{X}$ is a 5-dominating set of $C^{\prime}$.
Proof.
[…]
Finally, we clearly have $|Z_{X}|\leq|X|=\mu$. It hence follows from Proposition 2.2 that $tw(C)\leq tw(C^{\prime})=O(\sqrt{\mu})$. ∎
With Lemma C.19, we can now show using Theorem 2.3 that the above described distance property for minimum multicuts guarantees an efficient algorithm for planar instances.
Lemma C.21.
Let $(G,H)$ be a planar instance of Multicut such that for every minimum-weight multicut $S$ for $(G,H)$ and every $v\in V(G)$, we have $\textup{dist}_{G,S}(v,X)\leq 2$. Then an optimum solution for $(G,H)$ can be computed in $f(t)n^{O(\sqrt{\mu})}$.
Proof.
Let $C$ be an inclusionwise minimal, minimum weigh-multicut dual for $(G,H)$ and let $S=e_{G}(C)$. By Lemma 2.1, we have that $S$ is a minimum-weight multicut for $(G,H)$. It hence follows by assumption that $\textup{dist}_{G,S}(v,X)\leq 2$ holds for all $v\in V(G)$. It now follows from Lemma C.19 that $\textup{tw}(C)=O(\sqrt{\mu})$. The statement now follows from Theorem 2.3 for $F^{*}=\emptyset$. ∎
The main technical contribution for Lemma 3.4 is contained in the following result where the planar instance of Multicut mentioned above is constructed and its relationship to $(G,H)$ is described.
Lemma C.22.
Let $(G,H)$ be a connected $\textup{\sf planar}+\pi e$ instance of Multicut and let a complete maximum valid state $\mathcal{P}$ of $(G,H)$ with $\mathcal{P}_{\textup{thin}}=\emptyset$ be given. Then, in polynomial time, we can compute a weighting $G^{\prime}_{0}$ of $G_{0}$ and a graph $H^{\prime}$ with $V(H^{\prime})\subseteq V(G)$ and $|V(H^{\prime})|\leq\tau(\mathcal{P})$ such that there exists some $S_{\pi}\subseteq E_{\pi}$ with the following properties:
  1. 1.
    for every minimum multicut $S_{0}^{\prime}$ of $(G_{0}^{\prime},H^{\prime})$, we have that $S_{0}^{\prime}\cup S_{\pi}$ is a minimum multicut for $(G,H)$,
  2. 2.
    a minimum weight multicut for $(G^{\prime}_{0},H^{\prime})$ can be computed in $f(|V(H^{\prime})|)n^{O(\sqrt{\mu})}$.
Proof.
We first define $G^{\prime}_{0}$. Namely, for all $Y\in\mathcal{P}$ and every $e\in E(G_{0}[Y])$, we set the weight of $e$ to $\infty$. All remaining edges of $G^{\prime}_{0}$ are of weight $1$. We now define $H^{\prime}$. First, for every $Y\in\mathcal{P}$ and every component $Q$ of $G_{0}[Y]$, we choose a vertex $v_{Q}\in V(Q)$ and let $V(H^{\prime})$ contain $v_{Q}$. Next, for any distinct $Y_{1},Y_{2}\in\mathcal{P}$, any component $Q_{1}$ of $G_{0}[Y_{1}]$ and any component $Q_{2}$ of $G_{0}[Y_{2}]$, we let $V(H^{\prime})$ contain an edge linking $v_{Q_{1}}$ and $v_{Q_{2}}$. This finishes the description of $G^{\prime}_{0}$ and $H^{\prime}$. Clearly, $G^{\prime}_{0}$ and $H^{\prime}$ can be computed in polynomial time and $|V(H^{\prime})|\leq\tau(\mathcal{P})$.
Now, as $\mathcal{P}$ is valid, there exists a minimum multicut $S$ for $(G,H)$ that respects $\mathcal{P}$. Let $S_{\pi}=S\cap E_{\pi}$. Observe that $S_{0}$ is a multicut for $(G^{\prime}_{0},H^{\prime})$ as $S$ is a multicut for $(G,H)$ that respects $\mathcal{P}$ and by construction. Moreover, by the minimality of $S$, as $S$ respects $\mathcal{P}$ and by construction, we obtain that $S_{0}$ is of finite weight.
Claim C.23.
Let $S^{\prime}_{0}$ be a minimum multicut for $(G^{\prime}_{0},H^{\prime})$ and $S^{\prime}=S_{0}^{\prime}\cup S_{\pi}$. Then $S^{\prime}$ is a multicut for $(G,H)$.
Proof.
[…]
We are now ready to show the first property of the lemma.
Claim C.24.
Let $S^{\prime}_{0}$ be a minimum multicut for $(G^{\prime}_{0},H^{\prime})$ and $S^{\prime}=S_{0}^{\prime}\cup S_{\pi}$. Then $S^{\prime}$ is a minimum multicut for $(G,H)$.
Proof.
[…]
We still need to prove the second property of the lemma. To his end, we first give the following slight strengthening of Claim C.24.
Claim C.25.
Let $S^{\prime}_{0}$ be a minimum multicut for $(G^{\prime}_{0},H^{\prime})$ and $S^{\prime}=S_{0}^{\prime}\cup S_{\pi}$. Then $S^{\prime}$ is a minimum multicut for $(G,H)$ that respects $\mathcal{P}$.
Proof.
By Claim C.24, we only need to show that $S^{\prime}$ that respects $\mathcal{P}$. Let $Y_{1},Y_{2}\in\mathcal{P}$ with $Y_{1}\neq Y_{2}$ and let $y_{1}\in Y_{1}$ and $y_{2}\in Y_{2}$. We start by showing that $y_{1}$ and $y_{2}$ are in distinct components of $G^{\prime}_{0}\setminus S^{\prime}_{0}$. To this end, for $i\in[2]$, let $Q_{i}$ be the component of $G_{0}[Y_{i}]$ containing $y_{i}$. Observe that, as $S_{0}^{\prime}$ is finite and by construction, we have that $y_{i}$ and $v_{Q_{i}}$ are in the same connected component of $G^{\prime}_{0}\setminus S_{0}^{\prime}$. Next, by construction, we have $v_{Q_{1}}v_{Q_{2}}\in E(H^{\prime})$. As $S^{\prime}_{0}$ is a multicut for $(G_{0}^{\prime},H^{\prime})$, we obtain that $v_{Q_{1}}$ and $v_{Q_{2}}$ are in distinct components of $G_{0}^{\prime}\setminus S_{0}^{\prime}$. It follows that $y_{1}$ and $y_{2}$ are in distinct components of $G_{0}^{\prime}\setminus S_{0}^{\prime}$.
[…]
We now show the second statement of the lemma. To this end, let $C$ be an inclusionwise minimal minimum weight multicut dual for $(G^{\prime}_{0},H^{\prime})$. Let $S^{\prime}_{0}=e_{G^{\prime}_{0}}(C)$. By Lemma 2.1, we have that $S^{\prime}_{0}$ is a minimum weight multicut for $(G^{\prime}_{0},H^{\prime})$. Let $S^{\prime}=S^{\prime}_{0}\cup S_{\pi}$. It follows from Claim C.25 that $S^{\prime}$ is a minimum multicut for $(G,H)$ that respects $\mathcal{P}$, so for every $Y\in\mathcal{P}$, there exists a unique $\overline{Y}\in\mathcal{K}(G_{0}\setminus S^{\prime}_{0})$ with $Y\subseteq\overline{Y}$. It follows from Lemma 3.2 that $\textup{dist}_{G_{0}\setminus S^{\prime}_{0}}(v,X)\leq 1$ for all $v\in\overline{Y}$ and all $Y\in\mathcal{P}_{\textup{fat}}$. Next for every $Y\in\mathcal{P}_{\textup{f-n}}$, by definition, there exists some $Y^{\prime}\in\mathcal{P}_{\textup{fat}}$ such that $E(G_{0})$ contains an edge linking $Y$ and $Y^{\prime}$. Observe that this edge also links $\overline{Y}$ and $\overline{Y^{\prime}}$. It follows that $\textup{dist}_{G_{0}\setminus S^{\prime}_{0}}(v,X)\leq 2$ for all $v\in\overline{Y}$ and all $Y\in\mathcal{P}_{\textup{f-n}}$. As $\mathcal{P}_{\textup{thin}}=\emptyset$ by assumption and $\bigcup_{Y\in\mathcal{P}}\overline{Y}=V(G)$, we obtain that $\textup{dist}_{G_{0},S^{\prime}_{0}}(v,X)\leq 2$ holds for all $v\in V(G)$. It now follows from Lemma C.21 that a minimum multicut for $(G_{0}^{\prime},H^{\prime})$ can be computed in $f(|V(H^{\prime})|)n^{O(\sqrt{\mu})}$. Hence the second property of the lemma also holds. ∎
We are now ready to prove Lemma 3.4, which we restate here for convenience. With Lemma C.22 at hand, the proof follows easily. See 3.4
Proof.
By Lemma C.22, in polynomial time, we can compute a weighting $G_{0}^{\prime}$ of $G_{0}$ and a graph $H^{\prime}$ with $V(H^{\prime})\subseteq V(G)$ and $|V(H^{\prime})|=\tau(\mathcal{P})$ such that there exists a set $S_{\pi}\subseteq E_{\pi}$ such that the conclusion of Lemma C.22 holds.
We now describe our algorithm. Throughout the algorithm, we maintain a multicut $S$ for $(G,H)$, starting with $S=E(G)$. It follows from Lemma C.22 that there exists an algorithm that runs in $f(\pi,\tau(\mathcal{P}))n^{O(\sqrt{\mu})}$ that we can apply to $(G_{0}^{\prime},H^{\prime})$ and that returns a minimum multicut $S_{0}$ for $(G_{0}^{\prime},H^{\prime})$. Now consider some $S^{\prime}_{\pi}\subseteq E_{\pi}$ We check whether $S_{0}\cup S^{\prime}_{\pi}$ is a multicut for $(G,H)$ with $|S_{0}\cup S^{\prime}_{\pi}|<|S|$. If this is the case, then we replace $S$ by $S_{0}\cup S^{\prime}_{\pi}$. Observe that this last test can be executed in polynomial time by Proposition C.2, so for every $S^{\prime}_{\pi}$, the running time of this procedure is polynomial.
We now do this for all $S^{\prime}_{\pi}\subseteq E_{\pi}$. We return the assignment of $S$ after the last one of these iterations. Observe that, when $S^{\prime}_{\pi}=S_{\pi}$, then we replace $S$ by a minimum multicut for $(G,H)$. Hence the algorithm outputs a minimum multicut for $(G,H)$ indeed. Finally observe that there are only $f(\pi)$ choices for $S^{\prime}_{\pi}$. It follows that the total running time of our algorithm is $f(\pi,\tau(\mathcal{P}))n^{O(\sqrt{\mu})}$. ∎
C.6 Reducing to smaller instances
In this section, we combine the previously established subroutines to reduce a given $\textup{\sf planar}+\pi e$ instance of Multicut. More concretely, we prove Lemma 3.1. The rest of Section C.6 is dedicated to proving Lemma 3.1. In Lemma C.26, we show how an initial collection of states one of which is maximum valid can efficiently be made available. Next, in Lemma C.27, we show how to prove the result once a maximum valid state is available. Afterwards, Lemma 3.1 follows easily.
Lemma C.26.
Let $(G,H)$ be a connected $\textup{\sf planar}+\pi e$ instance of Multicut. Then, in $f(\pi,t)n^{O(1)}$, we can compute a collection $\mathbb{P}$ of $f(\pi,t)$ states such that at least one of them is maximum valid.
Proof.
We let $\mathbb{P}$ be the set of partitions of $W\cup T$. Clearly, $\mathbb{P}$ can be computed in $f(\pi,t)n^{O(1)}$. By definition, every $\mathcal{P}\in\mathbb{P}$ is a state. It hence suffices to prove that one of them is maximum valid. Let $S$ be a minimum multicut for $(G,H)$ that maximizes $|\mathcal{K}(G_{0}\setminus S_{0})|$ and let $\mathcal{P}$ be the partition of $W\cup T$ which is defined by $\mathcal{P}=\{Y\cap(W\cup T):Y\in\mathcal{K}(G_{0}\setminus S_{0})\}$. Observe that by Proposition C.9, we have that $Y\cap(W\cup T)\neq\emptyset$ for $Y\in\mathcal{K}(G_{0}\setminus S_{0})$ and hence $\mathcal{P}\in\mathbb{P}$. Then, clearly $\mathcal{K}(G_{0}\setminus S_{0})$ extends $\mathcal{P}$ and hence $S$ respects $\mathcal{P}$. It follows that $\mathcal{P}$ is valid. Further, we have $|\mathcal{P}|=|\mathcal{K}(G_{0}\setminus S_{0})|$, so $\mathcal{P}$ is maximum. ∎
The most difficult part of the recursive procedure for computing our minimum multicut is encoded in the following algorithm. It shows how the desired output can be obtained once a maximum valid state is available. The algorithm requires the construction of a search tree, which is slightly technical. However, the idea of the algorithm is conceptually not difficult now that all necessary subroutines are available.
Lemma C.27.
Let $(G,H)$ be a connected $\textup{\sf planar}+\pi e$ instance of Multicut and let a maximum valid state $\mathcal{P}$ be given. Then, in $f(\pi,t)n^{O(\sqrt{\mu})}$, we can run an algorithm that returns a set $S$ and a collection of extended subinstances $(S_{i},(G_{i},H_{i}))_{i\in[k]}$ of $(G,H)$ for some $k=f(\pi,t)$ such that for all $i\in[k]$, we have that $\pi(G_{i})+|V(H_{i})|<\pi(G)+|V(H)|$ holds and either $S$ is a minimum multicut for $(G,H)$ or there exists some $i\in[k]$ such that $(S_{i},(G_{i},H_{i}))$ is optimumbound.
Proof.
We create a rooted search tree each of whose nodes is associated to a state of $(G,H)$. First, we initialize the tree by its root which is associated to $\mathcal{P}$.
Now consider a leaf of the current tree that has not yet been considered after its creation and is associated to an incomplete state $\mathcal{P}^{0}$.
First suppose that $\mathcal{P}^{0}$ is not relevant. In that case, by Lemma 3.5, in polynomial time, we can compute a relevant state $\mathcal{P}^{1}$ with $|\mathcal{P}^{1}|=|\mathcal{P}^{0}|,\tau(\mathcal{P}^{1})\leq\tau(\mathcal{P}^{0}),\kappa(\mathcal{P}^{1})\geq\kappa(\mathcal{P}^{0})$ and such that $\mathcal{P}^{1}$ is maximum valid if $\mathcal{P}^{0}$ is. We create a single child of the leaf in consideration and associate it to $\mathcal{P}^{1}$.
Now suppose that $\mathcal{P}^{0}$ is relevant. By Lemma 3.6, there exists an algorithm that runs in polynomial time which we can apply to $\mathcal{P}^{0}$ and that returns a collection of $k$ states $(\mathcal{P}^{i})_{i\in[k]}$ for some $k\leq\pi+t$ such that $\kappa(\mathcal{P}^{i})>\kappa(\mathcal{P}^{0})$ and $\tau(\mathcal{P}^{i})\leq\tau(\mathcal{P}^{0})+1$ holds for $i\in[k]$ and such that, if $\mathcal{P}^{0}$ is maximum valid, then $\mathcal{P}^{i}$ is maximum valid for some $i\in[k]$. If the algorithm returns a collection of $k$ states $(\mathcal{P}^{i})_{i\in[k]}$ for some $k\leq\pi+t$ such that $\kappa(\mathcal{P}^{i})>\kappa(\mathcal{P}^{0})$ and $\tau(\mathcal{P}^{i})\leq\tau(\mathcal{P}^{0})+1$ for $i\in[k]$, then we add $k$ nodes to the rooted tree which are children of the node in consideration and associate each of them to the state $\mathcal{P}^{i}$ for some $i\in[k]$ in a way that $\mathcal{P}^{i}$ is associated to one of these nodes for $i\in[k]$.
We do this as long as there exists a leaf of the current tree which has not been considered after its creation. Let $U$ be the final tree at the end of this procedure. Observe that $\kappa(\mathcal{P})\geq 0$ by definition and, for every leaf of $U$ that is associated to a state $\mathcal{P}^{\prime}$, we have $\kappa(\mathcal{P}^{\prime})=f(\pi,t)$ as $\mathcal{P}^{\prime}$ is a state and by the definition of $\kappa$. Further, for every node that is associated to a state $\mathcal{P}^{0}$ and every child of this node that is associated to a state $\mathcal{P}^{1}$, we have $\kappa(\mathcal{P}^{1})\geq\kappa(\mathcal{P}^{0})$ by assumption. Moreover, if $\mathcal{P}^{0}$ is relevant, than the inequality is strict. As all children of nodes associated to an irrelevant state are relevant, it follows that the depth of $U$ is $f(\pi,t)$. In particular, $U$ is well-defined. Further observe that for every node that is associated to a state $\mathcal{P}^{0}$ and every child of this node that is associated to a state $\mathcal{P}^{1}$, we have $\tau(\mathcal{P}^{1})\leq\tau(\mathcal{P}^{0})+1$ by assumption. As the depth of $U$ is $f(\pi,t)$, we obtain that $\tau(\mathcal{P}^{0})=f(\pi,t)$ for every state $\mathcal{P}^{0}$ that is associated to a leaf of $U$. Further observe that every leaf of $U$ either corresponds to a state that is not maximum valid or to a complete state. Next observe that, as $\mathcal{P}$ is maximum valid and every node of $U$ that is associated to a maximum valid state has a child that is associated to a maximum valid state, we obtain that there exists a leaf of $U$ that is associated to a maximum valid complete state.
For the next part of the algorithm, we maintain a multicut $S$ for $(G,H)$, starting with $E(G)$. Now consider a leaf of $U$ which is associated to a complete state $\mathcal{P}^{0}$. Recall that $\tau(\mathcal{P}^{0})=f(\pi,t)$. It hence follows from Lemma 3.4 that there exists an algorithm we can apply to $(G,H)$ that runs in $f(\pi,t)n^{O(\sqrt{\mu})}$ and that returns a minimum multicut for $(G,H)$ if $\mathcal{P}^{0}$ is maximum valid and $\mathcal{P}^{0}_{\textup{thin}}=\emptyset$. We apply this algorithm to $(G,H)$ and if it returns a multicut $S^{\prime}$ for $(G,H)$ with $|S^{\prime}|<|S|$, then we replace $S$ by $S^{\prime}$. We do this for all leaves of $U$ which are associated to a complete state. The assignment of $S$ after the last of the iterations is the first part of our output.
For the last part of our algorithm, again consider a leaf of $U$ which is associated to a complete state $\mathcal{P}^{0}$. It follows from Lemma 3.3 that there exists an algorithm we can apply to $(G,H)$ that runs in polynomial time and that, in case that $\mathcal{P}^{0}$ is valid and contains a thin partition class, returns an optimumbound extended subinstance of $(G,H)$. If the algorithm returns an extended subinstance of $(G,H)$, we add it to the collection of extended subinstances we return. We do this for all leaves of $U$ which are associated to a complete state.
For the correctness of the algorithm, recall that there exists a leaf of $U$ that is associated to a maximum valid state $\mathcal{P}^{0}$. If $\mathcal{P}^{0}_{\textup{thin}}=\emptyset$, then it follows from the assumption on the algorithm and by construction that the returned set $S$ is a minimum multicut for $(G,H)$. If $\mathcal{P}^{0}_{\textup{thin}}\neq\emptyset$, then it follows from the assumption on the algorithm of Lemma 3.3 and by construction that the algorithm returns a collection of extended subinstances of $(G,H)$ containing an optimumbound one. Let $k$ be the number of leaves of $U$. Observe that the number of extended subinstances of $(G,H)$ that are returned is clearly bounded by $k$. Further observe that, as every node of $U$ has $f(\pi,t)$ children and the depth of $U$ is $f(\pi,t)$, we have that $k=f(\pi,t)$. Hence the size of the output is appropriate.
For the running time, observe that, as every node of $U$ has $f(\pi,t)$ children and the depth of $U$ is $f(\pi,t)$, we obtain that the number of nodes of $U$ is $f(\pi,t)$. It hence suffices to prove that the running time of the operations associated to every single node of $U$ is bounded by $f(\pi,t)n^{O(\sqrt{\mu})}$. For every node that is associated to an incomplete state, observe that by assumption, the algorithm for potentially computing the states associated to its children runs in polynomial time. Further, for all possible outputs, we can clearly check in polynomial time if it is a collection of states of the appropriate size. Hence the computation time associated to this node is polynomial. Now consider a node associated to a complete state. By assumption, the algorithm used for computing a potential multicut for $(G,H)$ associated to this node runs in $f(\pi,t)n^{O(\sqrt{\mu})}$. Further, we can check whether the output is a multicut in polynomial time by Proposition C.1. For the second part, the algorithm potentially computing an extended subinstance of $(G,H)$ runs in polynomial time by assumption. Finally, we can clearly check whether the output is an instance with the described properties in polynomial time. Hence the computation time associate to this node is also polynomial. This finishes the proof. ∎
We are now ready to prove Lemma 3.1, which we restate here for convenience. See 3.1
Proof.
We first use Lemma C.26 to create a collection of states $(\mathcal{P}^{i})_{i\in[k_{0}]}$ for some $k_{0}=f(\pi,t)$ such that one of them is maximum valid. Now, throughout the algorithm, we maintain a multicut $S$ for $(G,H)$ and a collection $\mathcal{I}$ of extended subinstances $(S^{\prime},(G^{\prime},H^{\prime}))$ of $(G,H)$ satisfying $\pi(G^{\prime})+|V(H^{\prime})|<\pi(G)+|V(H)|$. We initialize $S$ by $E(G)$ and $\mathcal{I}$ by $\emptyset$. It follows from Lemma C.27 that there exists an algorithm running in $f(\pi,t)n^{O(\sqrt{\mu})}$ that we can apply to $(G,H)$ and a state $\mathcal{P}$ and that, in case that $\mathcal{P}$ is maximum valid, returns a subset of $E(G)$ and a collection of extended subinstances of $(G,H)$ such that either the edge set is a minimum multicut for $(G,H)$ or one of the extended subinstances is optimumbound. We now apply this algorithm to $(G,H)$ and $\mathcal{P}^{i}$ for all $i\in[k_{0}]$. Whenever the output contains an edge set $S^{\prime}$ with $|S^{\prime}|<|S|$, we replace $S$ by $S^{\prime}$ and whenever the output contains a collection of extended subinstances of $(G,H)$, we add all these extended subinstances to $\mathcal{I}$. We make the algorithm return the assignments of $S$ and $\mathcal{I}$ after the last one of these iterations.
For the correctness of the algorithm, consider some $i\in[k_{0}]$ such that $\mathcal{P}^{i}$ is maximum valid. We obtain that when we apply the algorithm of Lemma C.27 to $(G,H)$ and $\mathcal{P}^{i}$, we obtain an output that contains either a minimum multicut $S^{\prime}$ for $(G,H)$ or an optimumbound extended subinstance of $(G,H)$. In the first case, we obtain that $S$ is assigned a minimum multicut for $(G,H)$ after this iteration and in the second case, we obtain that an obtimumbound extended subinstance of $(G,H)$ is added to $\mathcal{I}$ in this iteration. Finally observe, as $k_{0}=f(\pi,t)$ and for every $i\in[k_{0}]$, we add $f(\pi,t)$ extended subinstances of $(G,H)$ to $\mathcal{I}$, we have $|\mathcal{I}|=f(\pi,t)$ after the last one of these iterations.
For the running time of the algorithm, first observe that the algorithm of Lemma C.26 runs in $f(\pi,t)n^{O(1)}$ by assumption. Next observe that for every $i\in[k_{0}]$, the algorithm of Lemma C.27 runs in $f(\pi,t)n^{O(\sqrt{\mu})}$. Further, we can clearly check in polynomial time if the output is of the desired form. As $k_{0}=f(\pi,t)$, we obtain that the entire algorithm runs in $f(\pi,t)n^{O(\sqrt{\mu})}$. ∎
C.7 The final algorithm
We are now ready to prove Theorem 1.5, which we restate here for convenience. While this involves the careful creation of a search tree, again, the proof conecept is rather straight forward relying on Lemma 3.1.
See 1.5
Proof.
In a first phase of the algorithm, we create a rooted search tree such that each node $q$ of the search tree is associated to an extended subinstance $(S_{q}^{1},(G_{q},H_{q}))$ of $(G,H)$ and a multicut $S_{q}^{2}$ of $(G_{q},H_{q})$. We start by creating the root $r$ of this search tree and defining $S_{r}^{1}=\emptyset$ and $(G_{r},H_{r})=(G,H)$. In the following, while there exists a node $q$ that has not been considered after its creation and is associated to an instance $(S_{q}^{1},(G_{q},H_{q}))$ with $H_{q}$ not being edgeless, we compute $S_{q}^{2}$ and create its children together with the extended subinstance of $(G,H)$ associated to them.
Now consider a node $q$ that has not been considered after its creation. First suppose that $G_{q}$ is disconnected and let $G_{q}^{1},\ldots,G^{k}_{q}$ be the connected components of $G_{q}$ that satisfy $V(G^{i}_{q})\cap(W\cup T)\neq\emptyset$. Observe that $k\leq|W\cup T|=\pi+t$. We now add a collection of $k$ children $q_{1},\ldots,q_{k}$ of $q$ to the search tree. For $i\in[k]$, we set $S_{q_{i}}^{1}=\emptyset$ and $(G_{q_{i}},H_{q_{i}})=(G_{q}^{i},H[V(G_{q}^{i})])$. We further set $S_{q}^{2}=E(G_{q})$.
Now suppose that $G_{q}$ is connected. We first use the algorithm of Lemma C.3 to compute a set $E^{\prime}_{q}\subseteq E(G_{q})$ with $|E^{\prime}_{q}|=\pi(G_{q})$ such that $G_{q}\setminus E^{\prime}_{q}$ is planar. From now on, we may assume that $(G_{q},H_{q})$ is a $\textup{\sf planar}+\pi e$ instance. By Lemma 3.1 and as $(G_{q},H_{q})$ is a subinstance of $(G,H)$, in $f(\pi,t)n^{O(\sqrt{\mu})}$, we can run an algorithm that returns a set $S^{\prime}$ and a collection of extended subinstances $(S_{i},(G_{i},H_{i}))_{i\in[k]}$ of $(G,H)$ for some $k=f(\pi,t)$ such that for all $i\in[k]$, we have that $\pi(G_{i})+|V(H_{i})|<\pi(G_{q})+|V(H_{q})|$ holds and either $S^{\prime}$ is a minimum multicut for $(G_{q},H_{q})$ or there exists some $i\in[k]$ such that $(S_{i},(G_{i},H_{i}))$ is an optimumbound extended subinstance of $(G_{q},H_{q})$. We first set $S_{q}^{2}=S^{\prime}$. Then, for $i\in[k]$, we create a child $q_{i}$ of $q$ and set $(S_{q_{i}},(G_{q_{i}},H_{q_{i}}))=(S_{i},(G_{i},H_{i}))$.
We do this as long as there exists a node $q$ that has not been considered after its creation and is associated to an extended subinstance $(S_{q}^{1},(G_{q},H_{q}))$ of $(G,H)$ with $H_{q}$ not being edgeless. Finally, for all nodes $q$ which are associated to an extended subinstance $(S_{q}^{1},(G_{q},H_{q}))$ of $(G,H)$ with $H_{q}$ being edgeless, we set $S_{q}^{2}=\emptyset$. This finishes the description of the search tree, whose assignment in the end of the procedure we denote by $U$. This finishes the description of the first phase of the algorithm.
Observe that for every node $q$ of $U$ and every child $q^{\prime}$ of $q$ in $U$, we have $\pi(G_{q^{\prime}})+|V(H_{q^{\prime}})|\leq\pi(G_{q})+|V(H_{q})|$. Moreover, if $G_{q}$ is connected, then for every node $q$ of $U$ and every grandchild $q^{\prime}$ of $q$, we have that $G_{q^{\prime}}$ is connected and $\pi(G_{q^{\prime}})+|V(H_{q^{\prime}})|<\pi(G_{q})+|V(H_{q})|$ is satisfied. It follows that the depth of $U$ is bounded by $2(\pi+t)$. Further, by the assumption on the algorithm of Lemma 3.1 and construction, we have that every node of $U$ has $f(\pi,t)$ children. It follows that $U$ has $f(\pi,t)$ nodes in total. Now, for every node of $U$, we have that the running time associated to this node is $f(\pi,t)n^{O(\sqrt{\mu})}$ by the assumption on the algorithms of Lemmas 3.1 and C.3, and as the connected components of a graph can clearly be computed in polynomial time. It follows that the first phase of the algorithm runs in $f(\pi,t)n^{O(\sqrt{\mu})}$.
In the second phase of the algorithm, we compute a minimum multicut $S^{*}_{q}$ for $(G_{q},H_{q})$ for every node $q$ of $U$ by a bottom-up approach. First, for every leaf $q$ of $U$ which is associated to an extended subinstance $(S_{q}^{1},(G_{q},H_{q}))$ of $(G,H)$ with $H_{q}$ being edgeless, we set $S^{*}_{q}=\emptyset$. Clearly, we have that $\emptyset$ is a minimum multicut for $(G_{q},H_{q})$, so this assignment is correct. Now consider a node $q$ which is associated to an extended subinstance $(S_{q}^{1},(G_{q},H_{q}))$ of $(G,H)$ with $H_{q}$ not being edgeless and let $q_{1},\ldots,q_{k}$ be the children of $q$. Recursively, we may suppose that $S_{q_{i}}^{*}$ has already been computed for $i\in[k]$.
First suppose that $G_{q}$ is disconnected. In that case, let $S_{q}^{*}=\bigcup_{i\in[k]}S_{q_{i}}^{*}$, where $q_{1},\ldots,q_{k}$ are the children of $q$. It follows directly from the construction of $U$ that $S_{q}^{*}$ is indeed a minimum multicut for $(G_{q},H_{q})$.
Now suppose that $G_{q}$ is connected. We now let $\mathcal{S}_{q}$ contain $S_{q}^{2}$ and $S_{q_{i}}^{*}\cup S_{q_{i}}^{1}$ for $i\in[k]$. We now let $S_{q}^{*}$ be the smallest multicut for $(G_{q},H_{q})$ that is contained in $\mathcal{S}_{q}$. It follows from the assumption on the algorithm of Lemma 3.1 that $S_{q}^{*}$ is indeed a minimum multicut for $(G_{q},H_{q})$. We now recursively compute $(G_{q},H_{q})$ for all nodes $q$ of $U$. In the end, we have in particular computed $S_{r}^{*}$, which is a minimum multicut for $(G,H)$. We output $S_{r}^{*}$.
It is not difficult to see that for every node $q$ of $U$, we can compute $S_{q}^{*}$ in $f(\pi,t)n^{O(1)}$ once $S_{q^{\prime}}^{*}$ is computed for all children $q^{\prime}$ of $q$. As the number of nodes of $U$ is $f(\pi,t)$, we obtain that the second phase of the algorithm runs in $f(\pi,t)n^{O(1)}$. Hence the overall running time of the algorithm is $f(\pi,t)n^{O(\sqrt{\mu})}$. ∎
Appendix D Proof of Theorem 1.7
We here show how to conclude Theorem 1.7 from Theorem 4.2.
Proof.
By Proposition 4.1 and as the running time in Proposition 4.1 is dominated by the running time in Theorem 1.7, we may suppose that $(G,H)$ is given in the form of a crossing-planar instance with $\overline{cr}\leq 2cr$. Consider some $Z\subseteq E_{cr}$. Then we let $G_{Z}$ be obtained from $G$ by deleting $Z$ and assigning all edges in $E_{cr}\setminus Z$ infinite weight. Observe that $(G_{Z},H)$ is a normalized crossing-planar instance for every $Z\subseteq E_{cr}$. Using Theorem 4.2, we can hence compute a minimum weight multicut $S_{Z}$ for $(G_{Z},H)$ and check whether $S_{Z}\cup Z$ is a multicut for $(G,H)$. We let $S$ be the minimum weight multicut for $(G,H)$ found during this procedure. In order to see that $S$ is a minimum weight multicut for $(G,H)$, let $S^{*}$ be a minimum weight multicut for $(G,H)$ and let $Z^{*}=S^{*}\cap E_{cr}$. Then $S^{*}\setminus Z^{*}$ is a multicut for $(G_{Z^{*}},H)$ and hence a minimum weight multicut for $(G,H)$ is found when considering $Z=Z^{*}$. Finally, we execute the algorithm of Theorem 4.2 only $2^{\overline{cr}}$ times and hence the desired running time follows. ∎
Appendix E Proof of Theorem 4.2
We here show how to conclude Theorem 4.2 from Lemma 4.3.
Proof.
(of Theorem 4.2) By Lemma 4.3, there exists a graph $H^{*}\in\mathcal{H}$ such that for every minimum weight multicut $S_{0}$ for $(G^{\prime},H^{*})$, we have that $S_{0}$ is a minimum weight multicut for $(G,H)$ and for every subcubic inclusion-wise minimal minimum weight multicut dual $C$ for $(G^{\prime},H^{*})$, we have $\textup{tw}(C-(V_{F^{*}}^{C}\cup E_{F^{*}}^{C}))=O(\sqrt{\mu})$. Throughout the algorithm, we maintain a multicut $S$ for $(G,H)$, starting with $S=E(G)$.
Now let $H^{\prime}\in\mathcal{H}$. Observe that $(G^{\prime},H^{\prime})$ is a planar instance of Multicut. Hence, by Theorem 2.3, as $|F^{*}|\leq\overline{cr}$ and $|V(H^{\prime})|\leq t+2\overline{cr}$, there exists an algorithm we can apply to $(G^{\prime},H^{\prime})$ and that computes a minimum weight multicut for $(G^{\prime},H^{\prime})$ in $f(\overline{cr},t)n^{O(\sqrt{\mu})}$ if $\textup{tw}(C-(V_{F^{*}}^{C}\cup E_{F^{*}}^{C}))=O(\sqrt{\mu})$ holds for every subcubic inclusion-wise minimal minimum weight multicut dual $C$ for $(G^{\prime},H^{\prime})$. If the algorithm outputs a multicut $S_{H^{\prime}}$ for $(G,H)$ and the weight of $S_{H^{\prime}}$ is smaller than $w(S)$, then we replace $S$ by $S_{H^{\prime}}$, where $w$ is the weight function associated to $G$. Observe that, when applying this procedure for $H^{\prime}=H^{*}$, then $S$ is assigned a minimum weight multicut for $(G,H)$.
We do this for every $H^{\prime}\in\mathcal{H}$. Afterwards, we have that $S$ is assigned a minimum weight multicut for $(G,H)$. As there are only $f(\overline{cr},t)$ choices for $H^{\prime}$, we obtain that the total running time of the algorithm is $f(\overline{cr},t)n^{O(\sqrt{\mu})}$. ∎
Appendix F Proof of Lemma 4.3
The purpose of this appendix is to give the full proof of Lemma 4.3, which is a crucial ingredient in the proof of Theorem 1.7. We restate Lemma 4.3 here for convenience. Recall that $G^{\prime}$ is obtained from $G$ by deleting $E_{cr}$, $F^{*}$ is the set of faces of $G^{\prime}$ that in the correspodning embedding of $G$ contained an edge crossing and $\mathcal{H}$ is the set of all graphs on $V(H)\cup V(E_{cr})$. See 4.3
Proof.
Let $(B_{1},B_{2},I,X)$ be an extended biclique decomposition of $H$ with $|X|=\mu$. Further, let $S^{*}$ be a minimum weight multicut for $(G,H)$ such that $S^{*}$ minimizes the number of elements of $V(H)\cup V(E_{cr})$ that are in a component of $G\setminus S^{*}$ that also contains a terminal of $X$. In the following, we use $\overline{X}$ for the set of elements of $V(H)\cup V(E_{cr})$ that are in a component of $G\setminus S^{*}$ that also contains a terminal of $X$, we use $\overline{B_{1}}$ for the set of elements of $(V(H)\cup V(E_{cr}))\setminus\overline{X}$ that are in a component of $G\setminus S^{*}$ that also contains a terminal of $B_{1}$, and we use $\overline{B_{2}}$ for $(V(H)\cup V(E_{cr}))\setminus(\overline{X}\cup\overline{B_{1}})$. Observe that $(\overline{B_{1}},\overline{B_{2}},\overline{X})$ is a partition of $V(H)\cup V(E_{cr})$. We now define $H^{\prime}$ to be the complete multipartite graph such that $\overline{B_{1}}$ and $\overline{B_{2}}$ are partition classes and such that for every component of $G\setminus S^{*}$ that contains at least one terminal of $X$, all the elements of $\overline{X}$ contained in this component form a partition class.
This finishes the description of $H^{\prime}$. Observe that $H^{\prime}\in\mathcal{H}$. Further observe that $H$ is a subgraph of $H^{\prime}$. Indeed, consider some $t_{1},t_{2}\in V(H)$ with $t_{1}t_{2}\in E(H)$. If $\{t_{1},t_{2}\}\cap\overline{X}=\emptyset$, we have $\{t_{1},t_{2}\}\cap X=\emptyset$ and may hence suppose by symmetry that $t_{i}\in B_{i}$ for $i\in[2]$. It follows by construction and $S^{*}$ being a multicut for $(G,H)$ that $t_{i}\in\overline{B_{i}}$ for $i\in[2]$ yielding $t_{1}t_{2}\in E(H^{\prime})$ by construction. On the other hand, if one of $t_{1}$ and $t_{2}$ is contained in $\overline{X}$, then $t_{1}$ and $t_{2}$ are in distinct components of $G\setminus S^{*}$ as $S^{*}$ is a multicut for $(G,H)$ and so $t_{1}t_{2}\in E(H^{\prime})$ follows by construction.
In the following, we prove through two claims that every minimum weight multicut for $(G^{\prime},H^{\prime})$ is also a minimum weight multicut for $(G,H^{\prime})$. We first give the following result which will later be reused in the proof of the treewidth bound. See 4.4
Proof.
Let $t_{1},t_{2}\in V(H^{\prime})$ such that there exists a path $P$ from $t_{1}$ to $t_{2}$ in $G\setminus S^{\prime}$. Let $v_{0},\ldots,v_{k}$ be the vertices of $V(P)\cap V(H^{\prime})$ in the order they appear on $P$ when going from $t_{1}$ to $t_{2}$ with $v_{0}=t_{1}$ and $v_{k}=t_{2}$. As $H^{\prime}$ is a complete multipartite graph, it suffices to prove that $H^{\prime}$ does not contain an edge linking $v_{i-1}$ and $v_{i}$ for any $i\in[k]$.
Let $i\in[k]$. If $v_{i-1}v_{i}\in E_{cr}$, we obtain that $v_{i-1}v_{i}\notin S^{*}$ as the weight of $v_{i-1}v_{i}$ is infinite and the weight of $S^{*}$ is finite. It follows that $v_{i-1}$ and $v_{i}$ are in the same connected component of $G\setminus S^{*}$. By the choice of $\overline{X},\overline{B_{1}}$, and $\overline{B_{2}}$, we obtain that $v_{i-1}$ and $v_{i}$ are not linked by an edge of $E(H^{\prime})$. Otherwise, by construction, the subpath of $P$ linking $v_{i-1}$ and $v_{i}$ does not use an edge of $E_{cr}$ and hence $v_{i-1}$ and $v_{i}$ are in the same connected component of $G^{\prime}\setminus S^{\prime}$. As $S^{\prime}$ is a multicut for $(G^{\prime},H^{\prime})$, we obtain that $t_{1}$ and $t_{2}$ are not linked by an edge of $E(H^{\prime})$.
[…]
In order to prove the optimality, we further need the following result. See 4.5
Proof.
Let $t_{1},t_{2}\in V(H^{\prime})$ with $t_{1}t_{2}\in E(H^{\prime})$. If $\{t_{1},t_{2}\}\cap\overline{X}\neq\emptyset$, we obtain that $t_{1}$ and $t_{2}$ are in distinct components of $G\setminus S^{*}$ by the definition of $H^{\prime}$. As $G^{\prime}\setminus S^{*}$ is a subgraph of $G\setminus S^{*}$, we obtain that $t_{1}$ and $t_{2}$ are in distinct components of $G^{\prime}\setminus S^{*}$.
[…]
We are now ready to show that any minimum weight multicut $S$ for $(G^{\prime},H^{\prime})$ is a minimum weight multicut for $(G,H)$. First observe that $S$ is a multicut for $(G,H)$ by Claim 4.4 and as $H$ is a subgraph of $H^{\prime}$. Further, by Claim 4.5 and the minimality of $S$, we obtain $w(S)\leq w(S^{*})$ where $w$ is the weight function associated to $G$. By the minimality of $S^{*}$, the statement follows.
In the following, let $C$ be a subcubic, inclusion-wise minimal, minimum weight multicut dual for $(G^{\prime},H^{\prime})$. We will give the desired treewidth bound on $C$. Let $S^{\prime}=E_{G}(C)$ and observe that $S^{\prime}$ is a minimum weight multicut for $(G^{\prime},H^{\prime})$ by Lemma 2.1. It hence follows from the first part of the lemma that $S^{\prime}$ is also a minimum weight multicut for $(G,H)$. Further, we use $C^{\prime}$ for $C-F^{*}$.
We first need the following result. See 4.6
Proof.
Suppose otherwise. We will obtain a contradiction to the fact that $S^{*}$ is chosen so that the number of elements of $\overline{X}$ is minimized. Let $\overline{X}^{\prime}$ denote the set of terminals in $V(H^{\prime})$ that are contained in a component of $G\setminus S^{\prime}$ that also contains a terminal of $X$. As $f$ is a face of $C^{\prime}$, for every every $e\in E_{cr}$, either both or none of its endvertices are contained in $f$ and hence there is a collection of components of $G-S^{\prime}$ whose union is the set of vertices of $V(G)$ contained in $f$. By assumption, we obtain that $\overline{X}\setminus\overline{X}^{\prime}\neq\emptyset$.
Next consider some $x_{0}\in\overline{X}^{\prime}$. Let $K$ be the component of $G\setminus S^{\prime}$ containing $x_{0}$. By the definition of $\overline{X}^{\prime}$, we obtain that there exists some $x_{1}\in V(K)\cap X$. By Claim 4.4, we obtain that $S^{\prime}$ is a multicut for $(G,H^{\prime})$ and hence $x_{0}$ and $x_{1}$ are not linked by an edge of $E(H^{\prime})$. By the definition of $H^{\prime}$, we obtain that $x_{0}$ and $x_{1}$ are in the same component of $G\setminus S^{*}$. In particular, we have $x_{0}\in\overline{X}$. This yields $\overline{X}^{\prime}\subseteq\overline{X}$.
[…]
The next result is a crucial observation for the treewidth bound for $C$. See 4.7
Proof.
[…]
We are now ready to conclude the desired treewidth bound on $C^{\prime}$. Let $C^{\prime\prime}$ be a component of $C^{\prime}$. For every face $f$ of $C^{\prime\prime}$, we now introduce a vertex $z_{f}$ that shares an edge with every vertex that is incident to $f$. Let $C^{\prime\prime}_{0}$ be the resulting graph. Observe that $C^{\prime\prime}_{0}$ is planar. Now let $A\subseteq V(C^{\prime\prime}_{0})$ contain the vertex $z_{f}$ for all faces $f$ of $C^{\prime\prime}$ which contain a terminal of $X$. If $d_{C^{\prime\prime}}(v)\leq 2$ for all $v\in V(C^{\prime\prime})$, we clearly have $\textup{tw}(C^{\prime\prime})\leq 2$. Otherwise, it follows directly from Claims 4.6 and 4.7 and the fact that the faces of $C^{\prime\prime}$ correspond to the faces of $C^{\prime}$ that $A$ is a 3-dominating set of $C^{\prime\prime}_{0}$: if a face $f^{\prime}$ of $C^{\prime\prime}$ is adjacent to a face $f$ of $C^{\prime\prime}$ containing a vertex from $X$, then $z_{f^{\prime}}$ is at distance $2$ from $z_{f}$, and consequently every vertex incident to $f^{\prime}$ is at distance at most $3$ from $z_{f}$. By Proposition 2.2 and $|A|\leq|X|\leq\mu$, we obtain that $\textup{tw}(C^{\prime\prime}_{0})=O(\sqrt{\mu})$. As $C^{\prime\prime}$ is a subgraph of $C^{\prime\prime}_{0}$, we obtain that $\textup{tw}(C^{\prime\prime})=O(\sqrt{\mu})$. As $C^{\prime\prime}$ was chosen to be an arbitrary component of $C^{\prime}$, this yields $\textup{tw}(C^{\prime})=O(\sqrt{\mu})$. ∎